### v1

# A Concentration Bound for TD(0) with Function Approximation

Abstract
We derive a concentration bound of the type ‘for all $n\geq n_{0}$ for some $n_{0}$’ for TD(0) with linear function approximation. We work with online TD learning with samples from a single sample path of the underlying Markov chain. This makes our analysis significantly different from offline TD learning or TD learning with access to independent samples from the stationary distribution of the Markov chain. We treat TD(0) as a contractive stochastic approximation algorithm, with both martingale and Markov noises. Markov noise is handled using the Poisson equation and the lack of almost sure guarantees on boundedness of iterates is handled using the concept of relaxed concentration inequalities.
Keywords: TD(0), reinforcement learning, contractive stochastic approximation, concentration bound, temporal difference learning, all time bound
1 Introduction
TD(0) is one of the most popular reinforcement learning algorithms for policy evaluation (Tsitsiklis and Van Roy 1997). Given a fixed policy, the algorithm is an iterative method to obtain the value function for each state under the long-term discounted reward framework. To mitigate the issues of large state spaces, the value function is often approximated using a linear combination of feature vectors. This algorithm is referred to as TD(0) with linear function approximation. In this paper, we work with online TD(0) with a single sample path of the underlying Markov chain. Our goal in this paper is to obtain a bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
A bound of this form was published for TD(0) as a section in our paper titled “Concentration of Contractive Stochastic Approximation and Reinforcement Learning” (Chandak et al. 2022). This paper gave an all time bound for contractive stochastic approximation with Markov noise and applied the bound to asynchronous Q-learning and TD(0). Although the main theorem and its application to asynchronous Q-learning are correct, TD(0) does not satisfy a key assumption for the main theorem [^3] In Chandak et al. (2022), TD(0) does not satisfy assumption (6) which is required in the proof for Lemma 1 that shows almost sure boundedness of the iterates. The authors thank Zaiwei Chen for pointing this out. and hence the theorem was incorrectly applied to TD(0). We remove the need for that assumption in this version, giving a completely different proof specific to the TD(0) algorithm.
The previous paper required the iterates of the stochastic approximation iteration to be bounded by a constant with probability 1. This assumption is not known to be true for the iterates of online TD(0) with function approximation for a single sample path. In fact, a common method to alleviate this issue is to project the iterates back into a ball centered around the origin (Bhandari et al. 2018, Patil et al. 2023). The key difficulty caused by the lack of this assumption is in applying martingale inequalities which often require some restrictions on the increments of the martingale that are often not easy to verify. We do not work with the projection operator, and instead adapt relaxed concentration inequalities (Chung and Lu 2006, Section 8) for our problem. These bounds have an extra term given by the probability of increments going above a certain threshold (Tao and Vu 2015, Prop. 34).
While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$. In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$. Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al. (2021) (consider uniform sampling from dataset), Patil et al. (2023) (consider projection and tail averaging), and Chen et al. (2023), Dalal et al. (2018a) (consider i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$). Korda and LA (2015) give a high probability bound for the same online framework as us.
Going beyond TD(0), Bhandari et al. (2018), Chen et al. (2020, 2021), Wang et al. (2020), Dalal et al. (2018a, b), Li et al. (2020, 2021), Even-Dar and Mansour (2004), Prashanth et al. (2021), Qu and Wierman (2020), Sidford et al. (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for classes of algorithms. There is also a parallel body of work which seeks bound, either finite time or asymptotic, the difference between the value function under the learned policy and the optimal value function (Jin et al. 2018, Yang et al. 2020, Yang and Wang 2019), i.e., the ‘regret’. Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
Inspired by Chen et al. (2020), one of us considered stochastic approximation involving contractive maps and martingale noise, and derived such concentration bounds ‘from some $n_{0}$ on’ for this class of algorithms (Borkar 2021). This covered in particular synchronous Q-learning for discounted cost and some related schemes. In Chandak et al. (2022), we extended this work to cover ‘Markov noise’ (Meerkov 1972) in the stochastic approximation scheme, allowing us to give bounds for the asynchronous case. As mentioned before, this work assumed almost-sure boundedness of iterates, which is not satisfied by the TD(0) algorithm. We remove the need for this assumption in our current work. Other articles aiming at bounds of these forms, such as the one we provide, are found in Chandak et al. (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes.
To summarize, conventional concentration or sample complexity bounds yield bounds, e.g. mean square or ‘with high probability’ bounds, on the error from the target after $n$ iterations starting from time zero. The asymptotic regret bounds get an upper bound or lower bound or both on how some measure of cumulative error grows with time in an asymptotic sense. Our bounds differ from both. They are ‘all time bounds’ in the sense that they give a high probability bound for the iterates to remain in a prescribed small neighborhood of the target for all time from some time $n_{0}$ onwards. In fact the requirement ‘from some $n_{0}$ onwards’ is dictated by the fact that the decreasing stepsize needs to be sufficiently small from $n_{0}$ on. As for regret bounds, these are for cumulative error and are typically ‘almost sure’ or ‘in mean’ and asymptotic in nature unlike our bounds, which are from some time on, but with high probability.
The rest of the paper is structured as follows. Section 2 gives a background to the TD(0) algorithm, and states the main result. The result is proved in Section 3. A concluding section highlights some future directions. Appendix A states a martingale inequality used in our proof and Appendix B gives proofs for some technical lemmas which are used to prove the main theorem.
1.1 Notation
[…]
2 TD(0) - Background and Main Result
TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies (Tsitsiklis and Van Roy 1997). Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $S$. The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed. Assume that the chain is irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$. Let $D$ denote the $s\times s$ diagonal matrix whose $i$th diagonal entry is $\pi(i)$. Reward $r(i)$ is received when a transition from state $i$ takes place. The objective is to evaluate the long term discounted reward for each state given by the “value function”
$V(i)=E\left[\sum_{m=0}^{\infty}\gamma^{m}r(X_{m})|X_{0}=i\right],i\in S.$
Here $10<\gamma<$ 1 is a discount factor. The dynamic programming equation for evaluating the same is given by
$V(i)=r(i)+\gamma\sum_{j}p(j|i)V(j),\ i\in S,$
which can be written as the following vector equation
$V=r+\gamma PV$
for $r=[r(1),...,r(s)]^{T}$ and $P=[[p(j|i)]]_{i,j\in S}\in\mathcal{R}^{s\times s}$.
The state space can often be large ($s\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}:S\mapsto\mathcal{R},1\leq i\leq d$, with $s>>d\geq 1$. Also, let $\varphi(i)=[\phi_{1}(i),\ldots,\phi_{d}(i)]^{T}$. Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\phi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$. Here $x$ denotes the learnable weights for the linear function approximator. Since $\{\phi_{i}\}$ are linearly independent, $\Phi$ is full rank. Substituting this approximation into the dynamic programming equation above leads to
$\Phi x\approx r+\gamma P\Phi x.$
But the RHS may not belong to the range of $\Phi$. So we use the following fixed point equation:
$\Phi x=\Pi(r+\gamma P\Phi x)\coloneqq H(\Phi x),$ (1)
where $\Pi$ denotes the projection to Range($\Phi$) with respect to a suitable norm. It turns out to be convenient to take projection with respect to the weighted norm $\|x\|_{D}\coloneqq\sqrt{x^{T}Dx}=(\sum_{i}\pi(i)|x(i)|^{2})^{1/2}$ whereby the projection map is
$\Pi x\coloneqq\Phi(\Phi^{T}D\Phi)^{-1}\Phi^{T}Dx.$
The invertibility of $\Phi^{T}D\Phi$ is guaranteed by the fact that $\Phi$ is full rank. Finally, the TD(0) algorithm is given by the recursion
$x_{n+1}=x_{n}+a(n)\varphi(Y_{n})\Big{(}r(Y_{n})+\gamma\varphi(Y_{n+1})^{T}x_{n}-\varphi(Y_{n})^{T}x_{n}\Big{)}.$ (2)
Recall that $\varphi(i)^{T}\in\mathcal{R}^{d}$ for $i\in S$ denotes the $i$th row of $\Phi$ and $a(n)$ denotes the positive stepsize sequence. We now make the following two assumptions:
  1. 1.
    First we make an assumption on $\Phi$. For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$. Assume that
    $\lambda_{M}<\frac{\sqrt{2(1-\gamma)}}{(1+\gamma)}.$ (3)
    Since the feature vectors can be scaled without affecting the algorithm (the weights $x(i)$ get scaled accordingly), this assumption does not restrict the algorithm. Alternatively, the stepsize can be appropriately scaled as well, i.e., $a(n)=b(n)c$, where $b(n)$ acts as the effective stepsize and $c\varphi(\cdot)$ act as the effective feature vectors which satisfy the above assumption.
  2. 2.
    $\{a(n)\}$ is a sequence of non-negative stepsizes satisfying the conditions
    $a(n)\to 0,\ \sum_{n}a(n)=\infty.$
    and is assumed to be non-increasing, i.e., $a(n+1)\leq a(n)\ \forall\;n$. We also assume that $a(n)<1$ for all $n$. We further assume that $\frac{d_{1}}{n}\leq a(n)\leq d_{3}\left(\frac{1}{n}\right)^{d_{2}},\forall\ n$, where $d_{1}>0$ and $0<d_{2}\leq 1$. Larger values of $d_{1}$ and $d_{2}$ and smaller values of $d_{3}$ improve the main result presented below. The role this assumption plays in our bounds will become clear later.
    Observe that we do not require the classical square-summability condition in stochastic approximation, viz., $\sum_{n}a(n)^{2}<\infty$. This is because the contractive nature of our iterates (Lemma 1) gives us an additional handle on errors by putting less weight on past errors. A similar effect was observed in Chandak et al. (2022). The above assumptions on the stepsize sequence can be weakened so as to hold only after some $N>1$, without any changes in the analysis.
We next rearrange iteration (2) to separate the martingale noise and the ‘Markov noise’.
$x_{n+1}$ $=$ $x_{n}+a(n)\varphi(Y_{n})\Big{(}r(Y_{n})+\gamma\varphi(Y_{n+1})^{T}x_{n}-\varphi(Y_{n})^{T}x_{n}\Big{)}$
$=$ $x_{n}+a(n)\bigg{(}F(x_{n},Y_{n})-x_{n}+M_{n+1}x_{n}\bigg{)}$
$=$ $x_{n}+a(n)\left(\sum_{i\in S}\pi(i)F(x_{n},i)-x_{n}\right)+\underbrace{a(n)M_{n+1}x_{n}}_{\tau_{1}}+\underbrace{a(n)\left(F(x_{n},Y_{n})-\sum_{i\in S}\pi(i)F(x_{n},i)\right)}_{\tau_{2}}.$
where
$F(x,Y)=\varphi(Y)r(Y)+\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}x-\varphi(Y)\varphi(Y)^{T}x+x$
and
$M_{n+1}=\gamma\varphi(Y_{n})\Big{(}\varphi(Y_{n+1})^{T}-\sum_{j}p(j|Y_{n})\varphi(j)^{T}\Big{)}.$
Define the family of $\sigma$-fields $\mathcal{F}_{n}\coloneqq\sigma(x_{0},Y_{m},m\leq n),\ n\geq 0$. Then $\{M_{n}x_{n-1}\}$ is a martingale difference sequence, with respect to $\{\mathcal{F}_{n}\}$, i.e.,
$E[M_{n+1}x_{n}|\mathcal{F}_{n}]=\theta,a.s.\;\forall n,$
where $\theta$ denotes the zero vector. Then term $\tau_{1}$ denotes the error due to the martingale noise term and term $\tau_{2}$ denotes the error due to the ‘Markov noise’ $\{Y_{n}\}$.
The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. The proof for this lemma has been postponed to Appendix B.
Lemma 1.
For any $x,z\in\mathcal{R}^{d}$,
$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|\leq\alpha\|x-z\|,$
where
$\alpha=\sqrt{1-\min_{r\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
Moreover, $10<\alpha<$ 1 and hence the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction.
The Banach contraction mapping theorem implies that $\sum_{i}\pi(i)F(\cdot,i)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{i}\pi(i)F(x^{*},i)=x^{*}$. Then,
$\sum_{i}\pi(i)F(x^{*},i)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x^{*}=x^{*}$
$\implies(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi)x^{*}=\Phi^{T}D\Phi x^{*}$
$\implies(\Phi^{T}D\Phi)^{-1}(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi)x^{*}=x^{*}$
$\implies\Phi(\Phi^{T}D\Phi)^{-1}\Phi^{T}D(r+\gamma P\Phi)x^{*}=\Phi x^{*}$
$\implies H(\Phi x^{*})=\Phi x^{*}.$
So $\Phi x^{*}$ is the required fixed point of (1).
For $n\geq 0$, we further define:
[…]
$\beta_{k}(n)$ $=$ $\begin{cases}\frac{1}{k^{d_{2}-d_{1}}n^{d_{1}}},&\text{if}\ d_{1}\leq d_{2}\\ \frac{1}{n^{d_{2}}},&\text{otherwise}.\end{cases}$
Our main result is as follows:
Theorem 1.
There exist finite positive constants $c_{1},c_{2},c_{3}$ and $D$ such that for $0<\delta\leq 1,0<\epsilon\leq 1$, $n_{0}>0$ large enough to satisfy $\alpha+a(n_{0})c_{1}<1$, and $n\geq n_{0}$,
  1. (a)
    the inequality
    $\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n,$
    holds with probability exceeding
    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),0<\delta\leq C,$
    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),\delta>C.$
  2. (b)
    In particular,
    $\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall m\geq n_{0},$
    holds with probability exceeding
    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),0<\delta\leq C,$
    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),\delta>C.$
Here $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$.
The following are some remarks about the theorem and the proof that follows:
Remark 1.
The assumption that $\delta\leq 1$ and $\epsilon\leq 1$ is used in the proof for Lemma 3 and has been made only for simplicity. These can be taken as any positive values, with changes required only in the constants $C$ and $D$.
Remark 2.
The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$. This can be bounded using high probability bounds for TD(0) (e.g., Korda and LA (2015)) can be used as well. Alternatively, moment bounds (Srikant and Ying 2019, Bhandari et al. 2018, Chen et al. 2021) can be combined with Markov’s inequality to bound this term.
3 Proof of the Main Result
We present the proof of the main theorem in this section. The key martingale concentration inequality used in our proof is stated in Appendix A and proofs for the technical lemmas used in the proof are presented in Appendix B.
Proof.
Define $z_{n}$ for $n\geq n_{0}$ by
$z_{n+1}=z_{n}+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-z_{n}\right),$
where $z_{n_{0}}=x_{n_{0}}$. Note that $\|x_{n}-x^{*}\|\leq\|x_{n}-z_{n}\|+\|z_{n}-z^{*}\|$. To bound the second term, note that
$z_{n+1}-x^{*}$ $=(1-a(n))(z_{n}-x^{*})+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-x^{*}\right)$
$=(1-a(n))(z_{n}-x^{*})+a(n)\sum_{i}\pi(i)\left(F(z_{n},i)-F(x^{*},i)\right).$
The second equality follows from the fact that $x^{*}$ is a fixed point for $\sum_{i}\pi(i)F(\cdot,i)$. Then,
$\|z_{n+1}-x^{*}\|$ $\leq(1-a(n))\|z_{n}-x^{*}\|+a(n)\|\sum_{i\in S}\pi_{x_{n}}(i)(F(z_{n},i)-F(x^{*},i))\|$
$\leq(1-(1-\alpha)a(n))\|z_{n}-x^{*}\|.$
which finally gives us
$\|z_{n}-x^{*}\|\leq\psi(n,n_{0})\|x_{n_{0}}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(n-1)}\|x_{n_{0}}-x^{*}\|,$ (4)
where $\psi(n,m)\coloneqq\prod_{k=m}^{n-1}(1-(1-\alpha)a(k))$ if $n>m$ and $1$ otherwise. This also implies that for all $n\geq n_{0}$
$\|z_{n}\|\leq\|x_{n_{0}}-x^{*}\|+\|x^{*}\|.$ (5)
Next we give a probabilistic bound on the term $\|x_{n}-z_{n}\|$. Note that
$x_{n+1}-z_{n+1}=$ $(1-a(n))(x_{n}-z_{n})+a(n)M_{n+1}x_{n}$
$+a(n)\left(F(x_{n},Y_{n})-\sum_{i}\pi(i)F(z_{n},i)\right)$
$=(1-a(n))(x_{n}-z_{n})+\ a(n)M_{n+1}x_{n}$
$+\ a(n)\left(\sum_{i}\pi(i)(F(x_{n},i)-F(z_{n},i))\right)$
$+\ a(n)\left(F(x_{n},Y_{n})-\sum_{i}\pi(i)F(x_{n},i)\right).$
For $n,m\geq 0$, let $\chi(n,m)=\prod_{k=m}^{n}(1-a(k))$ if $n\geq m$ and $1$ otherwise. For some $n\geq n_{0}$, we iterate the above for $n_{0}\leq m\leq n$ to obtain
$x_{m+1}-z_{m+1}=$ $\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)M_{k+1}x_{k}$
$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)$
$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right).$ (6)
We first simplify the third term above. For simplicity, we define $F(x,Y)=F_{1}(Y)+F_{2}(Y)x+x,$ where
$F_{1}(Y)=\varphi(Y)r(Y)\in\mathcal{R}^{d}\;\;\;\textrm{and}\;\;\;F_{2}(Y)=\left(\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}-\varphi(Y)\varphi(Y)^{T}\right)\in\mathcal{R}^{d\times d}.$
We define $U:S\mapsto\mathcal{R}^{d}$ to be a solution of the Poisson equation:
$U(i)=F_{1}(i)-\sum_{j}\pi(j)F_{1}(j)+\sum_{j}p(j|i)U(j),\ i\in S.$ (7)
For $i_{0}\in S$, $\tau\coloneqq\min\{n>0:Y_{n}=i_{0}\}$ and $E_{i}[\cdots]=E[\cdots|Y_{0}=i]$, we know that
$U^{\prime}(i)=E_{i}\left[\sum_{m=0}^{\tau-1}(F_{1}(Y_{m})-\sum_{j\in\mathcal{S}}\pi(j)F_{1}(j))\right],\ i\in S,$ (8)
is one particular solution to the Poisson equation (see, e.g., Lemma 4.2 and Theorem 4.2 of Section VI.4, pp. 85-91, of Borkar (1991)). Thus $\|U^{\prime}(i)\|_{\infty}\leq 2\max_{i}\|F_{1}(i)\|_{\infty}E_{i}[\tau]$. For an irreducible Markov chain with a finite state space, $E_{i}[\tau]$ is finite for all $i$ and hence the solution $U^{\prime}(i)$ is bounded for all $i$. For each $\ell$, the Poisson equation specifies $U^{\ell}(\cdot)$ uniquely only up to an additive constant. Along with the additional constraint that $U(i_{0})=0$ for a prescribed $i_{0}\in S$, the system of equations given by (7) has a unique solution. Henceforth $U$ refers to the unique solution of the Poisson equation with $U(i_{0})=0$. Similarly, let $W:S\mapsto\mathcal{R}^{d\times d}$ be a solution of the Poisson equation:
$W(i)=F_{2}(i)-\sum_{j}\pi(j)F_{2}(j)+\sum_{j}p(j|i)W(j),\ i\in S,$ (9)
with the additional constraint that $W(i_{0})=0$ for a prescribed $i_{0}\in S$ as above.
The following lemma gives a simplification of the third term in (3), using the solutions of Poisson equation stated above. Before stating the lemma, we first define $x_{m}^{\prime}=\sup_{n_{0}\leq k\leq m}\|x_{m}-z_{m}\|$.
Lemma 2.
There exist positive constants $c_{1},c_{2}$ such that for all $n_{0}\leq m\leq n$,
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)$
$=\ \sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\widetilde{U}_{k+1}+\widetilde{W}_{k+1}x_{k}\right)+\mu_{m}(n_{0}),$
where
$\|\mu_{m}(n_{0})\|\leq a(n_{0})(c_{2}+c_{1}x_{m}^{\prime}+c_{1}\|x_{n_{0}}-x^{*}\|).$
Here $\widetilde{U}_{k+1}$ and $\widetilde{W}_{k+1}x_{k}$ are martingale difference sequences with respect to $\mathcal{F}_{k}$ where $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)$ and $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{j}p(j|Y_{k})W(j)$ for $k\geq n_{0}$ and and the zero vector, resp. the zero matrix, otherwise.
The proof appears in Appendix B. Returning to (3), we now have
$x_{m+1}-z_{m+1}$ $=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)$
$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)+\mu_{m}(n_{0})$
Now,
$\|x_{m+1}-z_{m+1}\|$ $\leq\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)\right\|$
$\;\;\;+\ \left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|$
$\;\;\;+\ a(n_{0})\left(c_{2}+c_{1}x_{m}^{\prime}+c_{1}\|x_{n_{0}}-x^{*}\|\right)$
$\leq\alpha\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\|x_{k}-z_{k}\|+a(n_{0})\left(c_{2}+c_{1}x_{m}^{\prime}+c_{1}\|x_{n_{0}}-x^{*}\|\right)$
$\;\;\;+\ \left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|.$ (10)
For any $0<k\leq m$,
$\chi(m,k)+\chi(m,k+1)a(k)=\chi(m,k+1),$
and hence
$\chi(m,n_{0})+\sum_{k={n_{0}}}^{m}\chi(m,k+1)a(k)=\chi(m,m+1)=1.$
This implies that
[…]
Using the definition of $x^{\prime}_{m}$, we have
$x^{\prime}_{m+1}\leq(\alpha+a(n_{0})c_{1})x^{\prime}_{m}+\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
Next we obtain a bound on the probability
$P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right),$
for some $\epsilon>0$ and $\delta>0$ (recall the assumption that $\alpha+a(n_{0})c_{1}<1$). From (4), recall that $\|z_{n}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(n-1))\|x_{n_{0}}-x^{*}\|\;a.s.,$ and hence,
$P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right)$
$>1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
To compensate for the lack of an almost sure bound on the iterates $\{x_{n}\}$, we adapt the proof method from Proposition 34 from Tao and Vu (2015) (see Section 8 of Chung and Lu (2006) for a detailed explanation). For this, we define $\xi=\{x_{0},Y_{k},k\geq 0\}$ and the ‘bad’ set $\mathbb{B}_{m}$ as
$\mathbb{B}_{m}=\left\{\xi\mid x_{m}^{\prime}(\xi)>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}(\xi)-x^{*}\|>\epsilon\right\}.$
Here the notation $x_{m}^{\prime}(\xi)$ and $x_{n_{0}}(\xi)$ highlights the dependence of $x_{m}^{\prime}$ and $x_{n_{0}}$ on the realizations of $x_{0}$ and $\{Y_{k}\}$. Analogous notation is used for other random variables. For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}$ and $\bar{z}_{k,n-1}(\xi)=z_{k}$ for all $k$. For $\xi\in\mathbb{B}_{n-1}$, we define $\bar{x}_{k,n-1}(\xi)=x^{*}$, and $\bar{z}_{k,n-1}(\xi)=x^{*}$ for all $k$. Also, define $\bar{x}_{n,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq n}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$. Now, adapting the proof of Proposition 34 from Tao and Vu (2015) for our case, we first note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies
$P(\bar{x}_{n,n-1}^{\prime}(\xi)\neq x_{n}^{\prime}(\xi))\leq P(\mathbb{B}_{n-1}).$
Henceforth we drop $\xi$ for ease of notation, rendering implicit the dependence of all random variables on $\xi$. Then,
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ $\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
The second inequality here follows from the observation that
$\left\{\bar{x}^{\prime}_{n,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcap\{\bar{x}^{\prime}_{n,n-1}=x^{\prime}_{n}\}\subseteq\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\},$
which implies that
$\left\{x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\subseteq\left\{\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcup\{\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\},$
which gives us the required inequality. The third inequality follows from union bound.
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$. Note that $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Using the definitions above,
$P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)=P\left(\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right).$
We can use (3) to obtain a bound on this probability. In the case where $\xi_{n-1}\in\mathbb{B}_{n-1}$, $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|=0$ and in the case where $\xi_{n-1}\notin\mathbb{B}_{n-1}$, we have $x_{k}=\bar{x}_{k,n-1}$. So (3) can be written as:
$\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$ $\leq(\alpha+a(n_{0})c_{1})\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}+a(n_{0})(c_{2}+c_{1}\epsilon)$
$\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|$
$\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)}{1-\alpha-a(n_{0})c_{1}}+\frac{\alpha+a(n_{0})c_{1}}{1-\alpha-a(n_{0})c_{1}}\delta$
$\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|$
When
$\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|\leq\delta,$
we have
$\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
Hence,
$P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
$\leq P\left(\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|>\delta\right).$
Let us denote the probability on the right side of the inequality as $p_{n}$. Then
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq p_{n}+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
Then repeating the same procedure using $\mathbb{B}_{n-2}$, we obtain
$P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq p_{n-1}+P\left(x_{n-2}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
Iterating this for $n\geq m\geq n_{0}+1$, we get
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq\sum_{m={n_{0}+1}}^{n}p_{m}+P(\|x_{n_{0}}-x^{*}\|>\epsilon).$
The following lemma, proved in Appendix B, gives a bound on the probabilities $p_{m}$:
Lemma 3.
There exist positive constants $c_{3}$ and $D$ such that for $0<\epsilon\leq 1,0<\delta\leq 1$,
$p_{m}\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)},0<\delta\leq C,$
$p_{m}\leq 2de^{-D\delta/\beta_{n_{0}}(m)},\delta>C,$
where $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$. Recall that $d$ here denotes the dimension of the iterates $\{x_{n}\}$.
This completes the proof for Theorem 1.
4 Conclusions
In conclusion, we note some future directions. The concept of relaxed martingale concentration inequalities to obtain bounds of the similar flavor for algorithms which suffer from similar issues. These include TD($\lambda$) and SSP Q Learning. Alternatively, similar bounds can be obtained for variants of temporal difference learning (Chen et al. 2021).
Appendix A Appendix A: A Martingale Inequality
Let $\{M_{n}\}$ be a real valued martingale difference sequence with respect to an increasing family of $\sigma$-fields $\{\mathcal{F}_{n}\}$. Assume that there exist $\varepsilon,C>0$ such that
[…]
Let $S_{n}:=\sum_{m=1}^{n}\zeta_{m,n}M_{m}$, where $\zeta_{m,n},\ m\leq n,$, for each $n$, are a.s. bounded $\{\mathcal{F}_{n}\}$-previsible random variables, i.e., $\zeta_{m,n}$ is $\mathcal{F}_{m-1}$-measurable $\forall\;m\geq 1$, and $|\zeta_{m,n}|\leq A_{m,n}$ a.s. for some constant $A_{m,n}$, $\forall\;m,n$. Suppose
$\sum_{m=1}^{n}A_{m,n}\leq\gamma_{1},\ \max_{1\leq m\leq n}A_{m,n}\leq\gamma_{2}\omega(n),$
for some $\gamma_{i},\omega(n)>0,\ i=1,2;n\geq 1$. Then we have:
Theorem 2.
There exists a constant $D>0$ depending on $\varepsilon,C,\gamma_{1},\gamma_{2}$ such that for $\epsilon>0$,
[…]
$2e^{-\frac{D\epsilon}{\omega(n)}},\ \ \mbox{otherwise.}$.
This is a variant of Theorem 1.1 of Liu and Watbled (2009). See Thoppe and Borkar (2019), Theorem A.1, pp. 21-23, for details.
Appendix B Appendix B: Technical Proofs
B.1 Proof of Lemma 1
Proof.
$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$ $=$ $\|\gamma\sum_{i}\pi(i)\varphi(i)\sum_{j}p(j|i)\varphi(j)^{T}(x-z)$ (13)
$\;\;\;\;\;\;\;-\sum_{i}\pi(i)\varphi(i)\varphi(i)^{T}(x-z)+(x-z)\|^{2}$
$=$ $\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)(x-z)\|^{2}$
$=$ $\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}$
$+(x-z)^{T}(x-z)$
$-2(x-z)^{T}\Phi^{T}D\Phi(x-z)$
$+(x-z)^{T}(\gamma\Phi^{T}DP\Phi+\gamma\Phi^{T}P^{T}D\Phi)(x-z).$
Now,
$(x-z)^{T}(\gamma\Phi^{T}DP\Phi+\gamma\Phi^{T}P^{T}D\Phi)(x-z)$ $=$ $(x-z)^{T}\gamma\Phi^{T}(DP+P^{T}D)\Phi(x-z)$ (14)
$=$ $\gamma\langle\Phi(x-z),P\Phi(x-z)\rangle_{D}$
$+\gamma\langle P\Phi(x-z),\Phi(x-z)\rangle_{D}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $2\gamma\|P\Phi(x-z)\|_{D}\|\Phi(x-z)\|_{D}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $2\gamma\|\Phi(x-z)\|_{D}^{2},$
and
$2(x-z)^{T}\Phi^{T}D\Phi(x-z)$ $=$ $2\langle\Phi(x-z),\Phi(x-z)\rangle_{D}$ (15)
$=$ $2\|\Phi(x-z)\|_{D}^{2}.$
Inequality (a) follows from the Cauchy-Schwarz inequality and (b) follows from the observation that $\|Px\|_{D}\leq\|x\|_{D}$, which can be proved as follows.
$\|Px\|^{2}_{D}=\sum_{i\in S}\pi(i)\left(\sum_{j\in S}p(j|i)x_{j}\right)^{2}\leq\sum_{i\in S}\pi(i)\sum_{j\in S}p(j|i)x_{j}^{2}=\sum_{j\in S}\pi(j)x_{j}^{2}=\|x\|^{2}_{D}.$
Here the inequality follows from Jensen’s inequality.
Combining (14) and (15) with (13) gives us:
$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$$\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}$ (16)
$+\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}.$
To analyze the last term in (16), we use the fact that the operator norm of a matrix defined as $\|M\|=\sup_{x\neq\theta}\frac{\|Mx\|}{\|x\|}$, using the Euclidean norm for vectors, is equal to the largest singular value of that matrix. Thus
$\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}$
$=\|\Phi^{T}\sqrt{D}(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
$\leq\lambda_{M}^{2}\|(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
$=\lambda_{M}^{2}\langle(\gamma P-I)\Phi(x-z),(\gamma P-I)\Phi(x-z)\rangle_{D}$
$=\lambda_{M}^{2}\|(I-\gamma P)\Phi(x-z)\|_{D}^{2}$
$\leq\lambda_{M}^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}.$ (17)
The last inequality follows from the triangle inequality. We now invoke assumption (3) and combine (B.1) with (16) as follows:
$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$$\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}+\lambda_{M}^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}$ (18)
$<$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}$
$+\left(\frac{\sqrt{2(1-\gamma)}}{1+\gamma}\right)^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}$
$=$ $\|x-z\|^{2}.$
This gives us the required contraction property with contraction factor $\alpha$ for which an explicit expression can be obtained, using the first inequality in (18), as
$\alpha=\sqrt{1-\min_{r\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
Note that as the columns of $\Phi$ are linearly independent, $x\neq\theta\implies\Phi x\neq\theta$ and hence $\frac{\|\Phi x\|_{D}}{\|x\|}>0$ when $x\neq\theta$. Along with assumption (3), this implies that $\alpha<1$. ∎
B.2 Proof of Lemma 2
Proof.
Using definitions of $U(\cdot)$ and $W(\cdot)$ we have,
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{j}p(j|Y_{k})U(j)\right)$ (19a)
$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{j}p(j|Y_{k})W(j)\right)x_{k}.$ (19b)
We first simplify (19a) as follows:
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{j}p(j|Y_{k})U(j)\right)$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)\right)$ (20a)
$\;\;\;+\sum_{k=n_{0}+1}^{m}\left((\chi(m,k+1)a(k)-\chi(m,k)a(k-1))U(Y_{k})\right)$ (20b)
$\;\;\;+\chi(m,n_{0}+1)a(n_{0})U(Y_{n_{0}})-\chi(m,m+1)a(m)U(Y_{m+1}).$ (20c)
For (20a), define $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)$ for $k\geq n_{0}$ and 00 otherwise. This is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
We define $U_{max}\coloneqq\max_{i\in S}\|U(i)\|$ and bound the norm of (20b) as follows:
$\left\|\sum_{k=n_{0}+1}^{m}\left((\chi(m,k+1)a(k)-\chi(m,k)a(k-1))U(Y_{k})\right)\right\|$
$\leq\left\|\sum_{k=n_{0}+1}^{m}((\chi(m,k+1)a(k)-\chi(m,k+1)a(k-1))U(Y_{k})\right\|$
$\;\;\;\;\;\;\;+\left\|\sum_{k=n_{0}+1}^{m}((\chi(m,k+1)a(k-1)-\chi(m,k)a(k-1))U(Y_{k})\right\|$
$\leq\sum_{k=n_{0}+1}^{m}((a(k-1)-a(k))\chi(m,k+1)U_{max})+\sum_{k=n_{0}+1}^{m}((\chi(m,k+1)-\chi(m,k))a(k-1)U_{max})$
$\leq\sum_{k=n_{0}+1}^{m}((a(k-1)-a(k))U_{max})+\sum_{k=n_{0}+1}^{m}((\chi(m,k+1)-\chi(m,k))a(n_{0})U_{max})$
$=(a(n_{0})-a(m))U_{max}+(\chi(m,m+1)-\chi(m,n_{0}+1))a(n_{0})U_{max}$
$\leq 2a(n_{0})U_{max}.$ (21)
The second and third inequality follow from $a(k-1)-a(k)\geq 0$ because $a(k)$ is a non-increasing sequence for $k>n_{0}$, and $\chi(m,k+1)-\chi(m,k)$ is positive because $1\geq\chi(m,k+1)\geq\chi(m,k)$ for $m,k>n_{0}$, as $a(k)<1$ for $k>n_{0}$. Note that the norm of (20c) is directly bounded by $2a(n_{0})U_{max}$.
Now we simplify (19b) as follows:
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{j}(p(j|Y_{k})W(j)\right)x_{k}$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k+1})-\sum_{j}(p(j|Y_{k})W(j)\right)x_{k}$ (22a)
$\;\;\;+\sum_{k=n_{0}+1}^{m}(\chi(m,k+1)a(k)-\chi(m,k)a(k-1))W(Y_{k})x_{k}$ (22b)
$\;\;\;+\sum_{k=n_{0}+1}^{m}\chi(m,k)a(k-1)W(Y_{k})(x_{k}-x_{k-1})$ (22c)
$\;\;\;+\chi(m,n_{0}+1)a(n_{0})W(Y_{n_{0}})x_{n_{0}}-\chi(m,m+1)a(m)W(Y_{m+1})x_{m}.$ (22d)
Similar to the sequence $\widetilde{U}_{k+1}$, for (22a), define $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{j}p(j|Y_{k})W(j)$ for $k\geq n_{0}$ and 00 otherwise. Note that $\widetilde{W}_{k+1}x_{k}$ is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
Define $W_{max}\coloneqq\max_{i\in S}\|W(i)\|$. Note that here $\|W(i)\|$ denotes the operator norm of a matrix, i.e., $\|W(i)\|=\sup_{x\neq\theta}\frac{\|W(i)x\|}{\|x\|}$, using the Euclidean norm for vectors. Similar to (20b), we bound the norm of (22b) as follows:
$\left\|\sum_{k=n_{0}+1}^{m}(\chi(m,k+1)a(k)-\chi(m,k)a(k-1))W(Y_{k})x_{k}\right\|$
$\leq\left\|\sum_{k=n_{0}+1}^{m}(\chi(m,k+1)a(k)-\chi(m,k)a(k-1))W(Y_{k})(x_{k}-z_{k})\right\|$
$\;\;\;+\left\|\sum_{k=n_{0}+1}^{m}(\chi(m,k+1)a(k)-\chi(m,k)a(k-1))W(Y_{k})z_{k}\right\|$
$\leq 2a(n_{0})W_{max}(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|).$
The last inequality here follows from the definition of $x_{m}^{\prime}=\sup_{n_{0}\leq k\leq m}\|x_{m}-z_{m}\|$ and from the bound on $\|z_{n}\|$ (5). For (22c), let us first bound $\|x_{k}-x_{k-1}\|$.
$\|x_{k}-x_{k-1}\|$ $=a(k)\left\|\varphi(Y_{k-1})\left(r(Y_{k-1})+\gamma\varphi(Y_{k})^{T}x_{k-1}-\varphi(Y_{k-1})^{T}x_{k-1}\right)\right\|$
$\leq a(k)\left(K_{1}+K_{2}\|x_{k-1}\|\right)$
$\leq a(n_{0})\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right)$
Before simplifying (22c), we first need to repeat an important simplification from our main proof. Note that for any $0<k\leq m$,
$\chi(m,k)+\chi(m,k+1)a(k)=\chi(m,k+1),$
and hence
$\chi(m,n_{0})+\sum_{k={n_{0}}}^{m}\chi(m,k+1)a(k)=\chi(m,m+1)=1.$
This implies that
[…]
We can finally bound the norm of (22c):
$\left\|\sum_{k=n_{0}+1}^{m}\chi(m,k)a(k-1)W(Y_{k})(x_{k}-x_{k-1})\right\|$
$\leq\sum_{k=n_{0}+1}^{m}\chi(m,k)a(k-1)\left\|W(Y_{k})(x_{k}-x_{k-1})\right\|$
$\leq\sum_{k=n_{0}+1}^{m}\chi(m,k)a(k-1)a(n_{0})W_{max}\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right)$
$\leq a(n_{0})W_{max}\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right).$
Finally the norm of (22d) can directly be bounded by
$\|\chi(m,n_{0}+1)a(n_{0})W(Y_{n_{0}})x_{n_{0}}-\chi(m,m+1)a(m)W(Y_{m+1})x_{m}\|$
$\leq 2a(n_{0})W_{max}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right).$
Combining the bounds above gives us
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\widetilde{U}_{k+1}+\widetilde{W}_{k+1}x_{k}\right)+\mu_{m}(n_{0}),$
where
$\mu_{m}(n_{0})\leq 4a(n_{0})U_{max}+a(n_{0})W_{max}\left(K_{1}+(4+K_{2})\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right).$
Define constants $c_{1}\coloneqq W_{max}(4+K_{2})$ and $c_{2}\coloneqq 4U_{max}+K_{1}W_{max}+c_{1}\|x^{*}\|$. This completes the proof for Lemma 2. ∎
B.3 Proof of Lemma 3
Proof.
We first note that, for $n_{0}\leq k\leq m$, $\|\bar{x}_{k,m}\|\leq\|\bar{x}_{k,m}-\bar{z}_{k,m}\|+\|\bar{z}_{k,m}\|\leq\bar{x}^{\prime}_{m}+\|\bar{z}_{k,m}\|$. By definition,
$\bar{x}^{\prime}_{m}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
Using (5), we have $\|\bar{z}_{k,m}\|\leq\epsilon+\|x^{*}\|$. Under the condition that $\epsilon\leq 1$ and $\delta\leq 1$, we have
$\|\bar{x}_{k,m}\|\leq 1+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}.$
Let $v^{(\ell)}$ denote the $\ell$th component of a vector $v$. Then
$\Gamma_{m}\coloneqq$ $\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)\right\|$
$\leq\sqrt{d}\max_{1\leq\ell\leq d}\left|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)^{(\ell)}\right|.$
Recall that $d$ here is the dimension of our iterates $\{x_{n}\}$. We apply Theorem 2 from Appendix A componentwise. For this, first note that
$\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)^{(\ell)}\leq c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right),$
where $c_{3}=\max\{M_{max}+2W_{max},2U_{max}\}$. In the theorem statement, let
$C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right),\zeta_{k,m}=\chi(m,k+1)a(k),\varepsilon=1,\gamma_{1}=1.$
Next, we choose suitable $\gamma_{2}$ and $\omega(m)$ such that $\max_{n_{0}\leq k\leq m}\zeta_{k,m}\leq\gamma_{2}\omega(m)$. For this, we use our assumption that $\frac{d_{1}}{n}\leq a(n)\leq d_{3}\left(\frac{1}{n}\right)^{d_{2}},\forall\ n\geq n_{0}$, to obtain:
$\chi(m,k+1)=\prod_{i=k+1}^{m}(1-a(i))\leq\exp\left(-\sum_{i=k+1}^{m}a(i)\right)\leq\exp\left(-\sum_{i=k+1}^{m}\frac{d_{1}}{i}\right)$
[…]
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\left(\frac{k+1}{k}\right)^{d_{1}}$
$\implies\max_{n_{0}\leq k\leq m}a(k)\chi(m,k+1)\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{k+1}{m}\right)^{d_{1}}\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{2k}{m}\right)^{d_{1}}.$
From the last inequality, $\gamma_{2}=d_{3}2^{d_{1}}$ and $\omega(m)=\beta_{n_{0}}(m)$ satisfy the required conditions. Then there exists a constant $D>0$, such that for $n_{0}<m$ and $\delta\in(0,C]$, we have
$P(\Gamma_{m}\geq\delta)\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)},$
and for $\delta>C$,
$P(\Gamma_{m}\geq\delta)\leq 2de^{-D\delta/\beta_{n_{0}}(m)}.$
The factor $d$ comes from the application of union bound to bound the maximum over all dimensions. ∎
Acknowledgements
The work of VSB was supported in part by S. S. Bhatnagar Fellowship from the Government of India.

### v2

@@
 In this paper, we work with online TD(0) with a single sample path of the underlying Markov chain.
-Our goal in this paper is to obtain a bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
+Our goal in this paper is to obtain a concentration bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
 A bound of this form was published for TD(0) as a section in our paper titled “Concentration of Contractive Stochastic Approximation and Reinforcement Learning” (Chandak et al.
@@
 samples of state - action - next state triplets $(s,a,s^{\prime})$).
-Korda and LA (2015) give a high probability bound for the same online framework as us.
+Korda and Prashanth (2015) give a high probability bound for the same online framework as us.
 Going beyond TD(0), Bhandari et al.
@@
 (2018a, b), Li et al.
-(2020, 2021), Even-Dar and Mansour (2004), Prashanth et al.
+(2020, 2023), Even-Dar and Mansour (2004), Prashanth et al.
 (2021), Qu and Wierman (2020), Sidford et al.
-(2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for classes of algorithms.
+(2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms.
 There is also a parallel body of work which seeks bound, either finite time or asymptotic, the difference between the value function under the learned policy and the optimal value function (Jin et al.
@@
 (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes.
+Another work which considers all-time bounds is Chen et al.
+(2023), but they only consider additive and multiplicative noise and not Markovian noise as considered in this paper.
 To summarize, conventional concentration or sample complexity bounds yield bounds, e.g.
@@
 Also, let $\varphi(i)=[\phi_{1}(i),\ldots,\phi_{d}(i)]^{T}$.
-Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\phi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$.
+Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\varphi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$.
 Here $x$ denotes the learnable weights for the linear function approximator.
@@
     First we make an assumption on $\Phi$.
-For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$.
+For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the square of the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$.
 Assume that
@@
 Alternatively, the stepsize can be appropriately scaled as well, i.e., $a(n)=b(n)c$, where $b(n)$ acts as the effective stepsize and $c\varphi(\cdot)$ act as the effective feature vectors which satisfy the above assumption.
+This assumption can be replaced with the following assumption on the basis vectors.
+    $\|\varphi(i)\|\leq\frac{\sqrt{2(1-\gamma)}}{1+\gamma}\;\forall i,$
+    i.e., the $\ell_{2}$ norm of each row of $\Phi$ is bounded by $\sqrt{2(1-\gamma)}/1+\gamma$. To see this, note that
+    $\frac{\|\Psi x\|_{2}}{\|x\|_{2}}=\frac{\|\Phi x\|_{D}}{\|x\|_{2}}\leq\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}.$
+    Now $\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}$ is the operator norm defined with $\ell_{2}$ norm for domain and $\ell_{\infty}$ for co-domain, which is equal to the maximum $\ell_{2}$ norm of a row.
+Hence
+    $\lambda_{M}=\max_{x\neq\theta}\frac{\|\Psi x\|_{2}}{\|x\|_{2}}\leq\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}=\max_{i}\|\varphi(i)\|_{2}.$
   2. 2.
@@
 Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$.
-The proof for this lemma has been postponed to Appendix B.
+Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor.
+Prior works such as Tsitsiklis and Van Roy (1997) show the contractive nature with respect to the weighted norm $\|\cdot\|_{D}$.
 Lemma 1.
@@
 where
-$\alpha=\sqrt{1-\min_{r\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
+$\alpha=\sqrt{1-\min_{x\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
 Moreover, $10<\alpha<$ 1 and hence the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction.
+The proof appears in Appendix B.
 The Banach contraction mapping theorem implies that $\sum_{i}\pi(i)F(\cdot,i)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{i}\pi(i)F(x^{*},i)=x^{*}$.
@@
     holds with probability exceeding
-    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),0<\delta\leq C,$
-    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),\delta>C.$
+    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
   2. (b)
@@
     holds with probability exceeding
-    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),0<\delta\leq C,$
-    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),\delta>C.$
-Here $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$.
+    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
 The following are some remarks about the theorem and the proof that follows:
@@
 The assumption that $\delta\leq 1$ and $\epsilon\leq 1$ is used in the proof for Lemma 3 and has been made only for simplicity.
-These can be taken as any positive values, with changes required only in the constants $C$ and $D$.
+These can be taken as any positive values, with changes required only in the constant $D$.
 Remark 2.
 The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$.
-This can be bounded using high probability bounds for TD(0) (e.g., Korda and LA (2015)) can be used as well.
+This can be bounded using high probability bounds for TD(0) (e.g., Korda and Prashanth (2015)).
 Alternatively, moment bounds (Srikant and Ying 2019, Bhandari et al.
@@
 2021) can be combined with Markov’s inequality to bound this term.
+Remark 3.
+For the special case of $a(n)=\frac{d_{1}}{n}$, and by combining our result with the finite time bound in Prashanth et al. (2021), we can rewrite the result in Theorem 1 in the following form.
+Corollary 1.
+Let $a(n)=d_{1}/n$ with sufficiently large $d_{1}$, and $n_{0}$ large enough to satisfy assumptions of Theorem 1. Then with probability at least $1-\varepsilon$, we have for all $m\geq n_{0}$ that
+$\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
+The proof for this corollary has been presented at the end of Appendix B.
+Chen et al.
+(2023) gives an all-time bound which matches our result at $n_{0}$, but their bound goes to zero as $m\uparrow\infty$.
+In our bound, the term $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ remains constant as $m$ is increased.
+Here $\delta$ can be modified to $\delta(m)$ (similar to the treatment in Corollary 1 in Chandak et al.
+(2022)).
+But the term $a(n_{0})(c_{2}+c_{1}\epsilon)$ arises from our treatment of Markov noise using Poisson equation.
+Note that Chen et al.
+(2023) do not consider the case of Markov noise, but only consider the case of additive and multiplicative noise (i.i.d.
+samples of state-action-next state triplets).
+We leave incorporating their ideas into our approach to get a bound decaying with $m$ for Markov noise as a future work.
 3 Proof of the Main Result
@@
 $z_{n+1}=z_{n}+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-z_{n}\right),$
-where $z_{n_{0}}=x_{n_{0}}$. Note that $\|x_{n}-x^{*}\|\leq\|x_{n}-z_{n}\|+\|z_{n}-z^{*}\|$. To bound the second term, note that
+where $z_{n_{0}}=x_{n_{0}}$. Note that $\|x_{n}-x^{*}\|\leq\|x_{n}-z_{n}\|+\|z_{n}-x^{*}\|$. To bound the second term, note that
 $z_{n+1}-x^{*}$ $=(1-a(n))(z_{n}-x^{*})+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-x^{*}\right)$
@@
 $+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right).$ (6)
-We first simplify the third term above. For simplicity, we define $F(x,Y)=F_{1}(Y)+F_{2}(Y)x+x,$ where
+Here we use the definition that $x_{n_{0}}=z_{n_{0}}$. We first simplify the third term above. For simplicity, we define $F(x,Y)=F_{1}(Y)+F_{2}(Y)x+x,$ where
 $F_{1}(Y)=\varphi(Y)r(Y)\in\mathcal{R}^{d}\;\;\;\textrm{and}\;\;\;F_{2}(Y)=\left(\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}-\varphi(Y)\varphi(Y)^{T}\right)\in\mathcal{R}^{d\times d}.$
@@
 Using the definition of $x^{\prime}_{m}$, we have
-$x^{\prime}_{m+1}\leq(\alpha+a(n_{0})c_{1})x^{\prime}_{m}+\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
+$x^{\prime}_{m+1}\leq$ $\;(\alpha+a(n_{0})c_{1})x^{\prime}_{m}+\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|$
+$+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
 Next we obtain a bound on the probability
@@
 From (4), recall that $\|z_{n}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(n-1))\|x_{n_{0}}-x^{*}\|\;a.s.,$ and hence,
+$\|x_{n_{0}}-x^{*}\|\leq\epsilon\implies\|z_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|.$
+Also recall that $\sup_{n_{0}\leq m\leq n}\|x_{m}-z_{m}\|=x_{n}^{\prime}$. Hence
+$\left\{\|x_{n_{0}}-x^{*}\|\leq\epsilon\right\}\bigcap\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}$
+$\subseteq\left\{\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right\}.$
+This implies the following relation between the probabilities of the two sets.
 $P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right)$
-$>1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
+$\geq 1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
 To compensate for the lack of an almost sure bound on the iterates $\{x_{n}\}$, we adapt the proof method from Proposition 34 from Tao and Vu (2015) (see Section 8 of Chung and Lu (2006) for a detailed explanation).
@@
 Analogous notation is used for other random variables.
-For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}$ and $\bar{z}_{k,n-1}(\xi)=z_{k}$ for all $k$.
+For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}(\xi)$ and $\bar{z}_{k,n-1}(\xi)=z_{k}(\xi)$ for all $k$.
 For $\xi\in\mathbb{B}_{n-1}$, we define $\bar{x}_{k,n-1}(\xi)=x^{*}$, and $\bar{z}_{k,n-1}(\xi)=x^{*}$ for all $k$.
-Also, define $\bar{x}_{n,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq n}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$.
+Also, define $\bar{x}_{m,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq m}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$.
 Now, adapting the proof of Proposition 34 from Tao and Vu (2015) for our case, we first note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies
@@
 Henceforth we drop $\xi$ for ease of notation, rendering implicit the dependence of all random variables on $\xi$. Then,
-$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ $\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
-$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
-$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
-The second inequality here follows from the observation that
+$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
+$\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\stackrel{{\scriptstyle(c)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
+$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
+Inequality (a) here follows from the observation that
 $\left\{\bar{x}^{\prime}_{n,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcap\{\bar{x}^{\prime}_{n,n-1}=x^{\prime}_{n}\}\subseteq\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\},$
@@
 $\left\{x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\subseteq\left\{\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcup\{\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\},$
-which gives us the required inequality. The third inequality follows from union bound.
-Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$.
-Note that $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$.
-Using the definitions above,
+which gives us the required inequality.
+Inequality (b) follows from union bound and inequality (c) follows from the observations that $\{\|x_{n_{0}}-x^{*}\|>\epsilon\}\subseteq\mathbb{B}_{n-1}$ and $\{\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\}\subseteq\mathbb{B}_{n-1}$.
+Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction.
+Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ for $n_{0}\leq m\leq n-1$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$.
+Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$.
+Then,
 $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)=P\left(\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right).$
@@
 where $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$. Recall that $d$ here denotes the dimension of the iterates $\{x_{n}\}$.
-This completes the proof for Theorem 1.
+This completes the proof for the first part of Theorem 1.
+Let $A_{n}$ be the set
+$\left\{\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right\}.$
+Then $\{A_{n}\}$ is a decreasing sequence of sets, i.e., $A_{n+1}\subseteq A_{n}$ for all $n\geq n_{0}$. Now let $A$ be the set
+$\left\{\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall m\geq n_{0}\right\}.$
+Then $A=\cap_{n=n_{0}}^{\infty}A_{n}$. Hence $P(A)=\lim_{n\uparrow\infty}P(A_{n})$. This completes the proof for Theorem 1.
 4 Conclusions
@@
 There exists a constant $D>0$ depending on $\varepsilon,C,\gamma_{1},\gamma_{2}$ such that for $\epsilon>0$,
-[…]
+$P\left(|S_{n}|>\epsilon\right)$$\leq$ $2e^{-\frac{D\epsilon^{2}}{\omega(n)}},\ \ \mbox{if}\ \epsilon\in\left(0,\frac{C\gamma_{1}}{\varepsilon}\right],$ (12)
 $2e^{-\frac{D\epsilon}{\omega(n)}},\ \ \mbox{otherwise.}$.
@@
 This gives us the required contraction property with contraction factor $\alpha$ for which an explicit expression can be obtained, using the first inequality in (18), as
-$\alpha=\sqrt{1-\min_{r\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
+$\alpha=\sqrt{1-\min_{x\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
 Note that as the columns of $\Phi$ are linearly independent, $x\neq\theta\implies\Phi x\neq\theta$ and hence $\frac{\|\Phi x\|_{D}}{\|x\|}>0$ when $x\neq\theta$.
+Also, note that $\min_{x\neq\theta}\frac{\|\Phi x\|_{D}}{\|x\|}=\min_{\|x\|=1}\|\Phi x\|_{D}$ and hence by extreme value theorem, we have that $\min_{\|x\|=1}\|\Phi x\|_{D}$ is attained and is greater than 00.
 Along with assumption (3), this implies that $\alpha<1$.
@@
 $P(\Gamma_{m}\geq\delta)\leq 2de^{-D\delta/\beta_{n_{0}}(m)}.$
-The factor $d$ comes from the application of union bound to bound the maximum over all dimensions. ∎
+The factor $d$ comes from the application of union bound to bound the maximum over all dimensions.
+Under the assumption that $\delta\leq 1$, we have that $e^{-D\delta^{2}/\beta_{n_{0}}(m)}\geq e^{-D\delta/\beta_{n_{0}}(m)}$ and hence $P(\Gamma_{m}\geq\delta)\leq 2de^{D\delta^{2}/\beta_{n_{0}}(m)}$.
+∎
+B.4 Proof of Corollary 1
+To show Corollary 1, we first obtain values of $\delta$ and $\epsilon$ such that the probability in Theorem 1 is $1-\varepsilon$.
+We use $\mathfrak{s}_{i},i=1,2,\ldots$ to denote different constants throughout this proof.
+For $a(n)=d_{1}/n$ with a sufficiently large $d_{1}$, we have $\beta_{n_{0}}(m)\leq 1/m$.
+This implies that
+$\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}m)\leq\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0}).$
+Let $\varepsilon/(4d)=\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0})$, which gives us $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon)$ for appropriate constants $\mathfrak{s}_{2}$ and $\mathfrak{s}_{3}$.
+This choice of $\delta$ gives us
+$22d\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\varepsilon/$ 2 .
+Using Theorem 1 from Korda and Prashanth (2015), we have $P(\|x_{n_{0}}-x^{*}\|\leq\epsilon)>1-\varepsilon/2$, where $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$.
+So, choosing $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$ gives us
+$P(\|x_{n_{0}}-x^{*}\|>\epsilon)\leq\varepsilon/2.$
+Next we substitute the values $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon)$ and $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$ in our bound from Theorem 1.
+Note that $\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon\leq\epsilon$.
+Since $a(n)=d_{1}/n$, we have $a(n_{0})c_{2}=\mathfrak{s}_{6}n_{0}^{-1}$ and $a(n_{0})c_{1}\epsilon=(\mathfrak{s}_{7}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{8})n_{0}^{-3/2}$.
+Combining these terms, we get the following.
+$e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ $\leq\mathfrak{s}_{9}n_{0}^{-1/2}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{10}n_{0}^{-1/2}$
+$=\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
+Hence we get that $\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right)$ holds for all $m\geq n_{0}$ with probability at least $1-\varepsilon$, completing the proof for Corollary 1.
 Acknowledgements

### v3

@@
 samples of state - action - next state triplets $(s,a,s^{\prime})$).
-Korda and Prashanth (2015) give a high probability bound for the same online framework as us.
 Going beyond TD(0), Bhandari et al.
@@
 (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms.
-There is also a parallel body of work which seeks bound, either finite time or asymptotic, the difference between the value function under the learned policy and the optimal value function (Jin et al.
+There is also a parallel body of work which seeks bound, either finite time or asymptotic, on the difference between the value function under the learned policy and the optimal value function (Jin et al.
 2018, Yang et al.
@@
 The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed.
-Assume that the chain is irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$.
+Assume that the chain is aperiodic irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$.
 Let $D$ denote the $s\times s$ diagonal matrix whose $i$th diagonal entry is $\pi(i)$.
@@
 We also assume that $a(n)<1$ for all $n$.
-We further assume that $\frac{d_{1}}{n}\leq a(n)\leq d_{3}\left(\frac{1}{n}\right)^{d_{2}},\forall\ n$, where $d_{1}>0$ and $0<d_{2}\leq 1$.
+We further assume that $\frac{d_{1}}{n+1}\leq a(n)\leq d_{3}\left(\frac{1}{n+1}\right)^{d_{2}},\forall\ n$, where $d_{1}>0$ and $0<d_{2}\leq 1$.
 Larger values of $d_{1}$ and $d_{2}$ and smaller values of $d_{3}$ improve the main result presented below.
@@
 Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor.
-Prior works such as Tsitsiklis and Van Roy (1997) show the contractive nature with respect to the weighted norm $\|\cdot\|_{D}$.
 Lemma 1.
@@
 The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$.
-This can be bounded using high probability bounds for TD(0) (e.g., Korda and Prashanth (2015)).
-Alternatively, moment bounds (Srikant and Ying 2019, Bhandari et al.
+This can be bounded by combining moment bounds (Srikant and Ying 2019, Bhandari et al.
 2018, Chen et al.
-2021) can be combined with Markov’s inequality to bound this term.
+2021) with Markov’s inequality.
 Remark 3.
-For the special case of $a(n)=\frac{d_{1}}{n}$, and by combining our result with the finite time bound in Prashanth et al. (2021), we can rewrite the result in Theorem 1 in the following form.
-Corollary 1.
-Let $a(n)=d_{1}/n$ with sufficiently large $d_{1}$, and $n_{0}$ large enough to satisfy assumptions of Theorem 1. Then with probability at least $1-\varepsilon$, we have for all $m\geq n_{0}$ that
-$\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
-The proof for this corollary has been presented at the end of Appendix B.
-Chen et al.
-(2023) gives an all-time bound which matches our result at $n_{0}$, but their bound goes to zero as $m\uparrow\infty$.
+In Chen et al.
+(2025), an all-time bound is obtained which goes to zero as $m\uparrow\infty$.
 In our bound, the term $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ remains constant as $m$ is increased.
@@
 We leave incorporating their ideas into our approach to get a bound decaying with $m$ for Markov noise as a future work.
+Remark 4.
+For the special case of $a(n)=\frac{d_{1}}{n+1}$, we combine our result with Theorem 2.1 from Chen et al. (2021), a mean square error bound, to get the following corollary.
+Corollary 1.
+Let $a(n)=d_{1}/(n+1)$ with sufficiently large $d_{1}$.
+Let $n_{0}$ be large enough to satisfy assumptions of Theorem 1 and (Chen et al.
+2021, Theorem 2.1).
+Then with probability at least $1-\varepsilon_{1}-\varepsilon_{2}$, we have for all $m\geq n_{0}$ that
+$\|x_{m}-x^{*}\|=\mathcal{O}\left(\frac{1}{\sqrt{n_{0}}}\log^{1/2}\left(\frac{1}{\varepsilon_{1}}\right)+\sqrt{\frac{\log(n_{0})}{n_{0}}}\frac{1}{\sqrt{\varepsilon_{2}}}\left(\frac{n_{0}}{m}+\frac{1}{n_{0}}\right)\right).$
+The proof for this corollary has been presented at the end of Appendix B.
+The first term here corresponds to the term $\delta$ in Theorem 1.
+This term has a $\sqrt{1/n_{0}}$ decay rate and an exponentially small tail.
+The second term is the contribution of the initial condition at $n_{0}$.
+We have a polynomial tail in this case, but the dependence on $m$ and $n_{0}$ is stronger, as the term $(\sqrt{\log(n_{0})}/\sqrt{n_{0}})\times(n_{0}/m)$ decays with $m$ and the other term decays as $\log^{1/2}(n_{0})n_{0}^{-3/2}$.
 3 Proof of the Main Result
@@
 The key martingale concentration inequality used in our proof is stated in Appendix A and proofs for the technical lemmas used in the proof are presented in Appendix B.
-Proof.
+Proof of Theorem 1..
 Define $z_{n}$ for $n\geq n_{0}$ by
@@
 Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction.
-Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ for $n_{0}\leq m\leq n-1$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$.
+Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$.
 Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$.
@@
 $\left\{\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall m\geq n_{0}\right\}.$
-Then $A=\cap_{n=n_{0}}^{\infty}A_{n}$. Hence $P(A)=\lim_{n\uparrow\infty}P(A_{n})$. This completes the proof for Theorem 1.
+Then $A=\cap_{n=n_{0}}^{\infty}A_{n}$. Hence $P(A)=\lim_{n\uparrow\infty}P(A_{n})$. This completes the proof for Theorem 1. ∎
 4 Conclusions
 In conclusion, we note some future directions.
-The concept of relaxed martingale concentration inequalities to obtain bounds of the similar flavor for algorithms which suffer from similar issues.
+The concept of relaxed martingale concentration inequalities can be used to obtain bounds of the similar flavor for algorithms which suffer from similar issues.
 These include TD($\lambda$) and SSP Q Learning.
@@
 2021).
+Another direction could be to improve the bounds in this paper to get an exponentially small tail for Markovian stochastic approximation.
 Appendix A Appendix A: A Martingale Inequality
@@
 B.4 Proof of Corollary 1
-To show Corollary 1, we first obtain values of $\delta$ and $\epsilon$ such that the probability in Theorem 1 is $1-\varepsilon$.
+To show Corollary 1, we first obtain values of $\delta$ and $\epsilon$ such that the probability in Theorem 1 is $1-\varepsilon_{1}-\varepsilon_{2}$.
 We use $\mathfrak{s}_{i},i=1,2,\ldots$ to denote different constants throughout this proof.
-For $a(n)=d_{1}/n$ with a sufficiently large $d_{1}$, we have $\beta_{n_{0}}(m)\leq 1/m$.
+For $a(n)=d_{1}/(n+1)$ with a sufficiently large $d_{1}$, we have $\beta_{n_{0}}(m)\leq 1/m$.
 This implies that
 $\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}m)\leq\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0}).$
-Let $\varepsilon/(4d)=\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0})$, which gives us $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon)$ for appropriate constants $\mathfrak{s}_{2}$ and $\mathfrak{s}_{3}$.
+Let $\varepsilon_{1}/(2d)=\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0})$, which gives us $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})$ for appropriate constants $\mathfrak{s}_{2}$ and $\mathfrak{s}_{3}$.
 This choice of $\delta$ gives us
-$22d\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\varepsilon/$ 2 .
-Using Theorem 1 from Korda and Prashanth (2015), we have $P(\|x_{n_{0}}-x^{*}\|\leq\epsilon)>1-\varepsilon/2$, where $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$.
-So, choosing $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$ gives us
-$P(\|x_{n_{0}}-x^{*}\|>\epsilon)\leq\varepsilon/2.$
-Next we substitute the values $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon)$ and $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$ in our bound from Theorem 1.
-Note that $\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon\leq\epsilon$.
-Since $a(n)=d_{1}/n$, we have $a(n_{0})c_{2}=\mathfrak{s}_{6}n_{0}^{-1}$ and $a(n_{0})c_{1}\epsilon=(\mathfrak{s}_{7}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{8})n_{0}^{-3/2}$.
-Combining these terms, we get the following.
-$e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ $\leq\mathfrak{s}_{9}n_{0}^{-1/2}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{10}n_{0}^{-1/2}$
-$=\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
-Hence we get that $\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right)$ holds for all $m\geq n_{0}$ with probability at least $1-\varepsilon$, completing the proof for Corollary 1.
+$2d\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\varepsilon_{1}.$
+Let $\epsilon=\sqrt{E[\|x_{n_{0}}-x^{*}\|^{2}]}/\sqrt{\varepsilon_{2}}$, which implies that
+$P(\|x_{n_{0}}-x^{*}\|>\epsilon)=P\left(\|x_{n_{0}}-x^{*}\|^{2}>\frac{E[\|x_{n_{0}}-x^{*}\|^{2}]}{\varepsilon_{2}}\right)\leq\varepsilon_{2}.$
+Here the last inequality follows from Markov’s inequality.
+Being linear contractive SA with an aperiodic irreducible Markov chain, our formulation satisfies the assumptions for Theorem 2.1 from Chen et al.
+(2021).
+To apply their result, we note the corresponding mapping between constants: the norm $\|\cdot\|_{c}$ is the Euclidean norm in our case, $h$ is $1$, $\varphi_{2}$ is $1-\alpha$ in our case and their $\alpha$ is $d_{1}$ in our case.
+For $d_{1}>1/(1-\alpha)$ and $n_{0}$ sufficiently large to satisfy the condition for Theorem 2.1 (2), we can use their result Theorem 2.1 (2)(b)(iii) to obtain the following mean square bound.
+$E\left[\|x_{n_{0}}-x^{*}\|^{2}\right]$ $\leq\mathfrak{s}_{4}\left(\frac{1}{n_{0}+1}\right)^{(1-\alpha)d_{1}}+\mathfrak{s}_{5}\frac{\log(n_{0}+1)}{n_{0}+1}\leq\mathfrak{s}_{6}\frac{\log(n_{0}+1)}{n_{0}+1}.$
+Substituting the values of $\delta$ and $\epsilon$ in our bound, we get with probability greater than $1-\varepsilon_{1}-\varepsilon_{2}$,
+$\|x_{m}-x^{*}\|$ $\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\frac{\sqrt{E[\|x_{n_{0}}-x^{*}\|^{2}]}}{\sqrt{\varepsilon_{2}}}$
+$\;\;+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\sqrt{E[\|x_{n_{0}}-x^{*}\|^{2}]}}{\sqrt{\varepsilon_{2}}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right)$
+$\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}$
+$+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right)$
+for all $m\geq n_{0}$. Now
+$\exp(-(1-\alpha)b_{n_{0}}(m-1))$
+$\leq\exp\left(-\sum_{i=n_{0}}^{m-1}(1-\alpha)a(i)\right)\leq\exp\left(-\sum_{i=n_{0}}^{m-1}\frac{(1-\alpha)d_{1}}{i+1}\right)$
+[…]
+$=\left(\frac{n_{0}+1}{m+1}\right)^{d_{1}(1-\alpha)}\leq\frac{n_{0}+1}{m+1}.$
+Here the final inequality follows from the assumption that $(1-\alpha)d_{1}>1$.
+Hence we get that for sufficiently large $n_{0}$, the following holds with probability $1-\varepsilon_{1}-\varepsilon_{2}$ for all $m\geq n_{0}$.
+$\|x_{m}-x^{*}\|=\mathcal{O}\left(\frac{1}{\sqrt{n_{0}}}\log^{1/2}\left(\frac{1}{\varepsilon_{1}}\right)+\sqrt{\frac{\log(n_{0})}{n_{0}}}\frac{1}{\sqrt{\varepsilon_{2}}}\left(\frac{n_{0}}{m}+\frac{1}{n_{0}}\right)\right).$
 Acknowledgements

### v4

@@
 Abstract
-We derive a concentration bound of the type ‘for all $n\geq n_{0}$ for some $n_{0}$’ for TD(0) with linear function approximation.
+We derive uniform all-time concentration bound of the type ‘for all $n\geq n_{0}$ for some $n_{0}$’ for TD(0) with linear function approximation.
 We work with online TD learning with samples from a single sample path of the underlying Markov chain.
@@
 1 Introduction
-TD(0) is one of the most popular reinforcement learning algorithms for policy evaluation (Tsitsiklis and Van Roy 1997).
+TD(0) is one of the most popular Reinforcement Learning (RL) algorithms for policy evaluation (Tsitsiklis and Van Roy 1997).
 Given a fixed policy, the algorithm is an iterative method to obtain the value function for each state under the long-term discounted reward framework.
@@
 2022).
-This paper gave an all time bound for contractive stochastic approximation with Markov noise and applied the bound to asynchronous Q-learning and TD(0).
+This paper established an all-time bound for contractive stochastic approximation with Markov noise and applied the bound to asynchronous Q-learning and TD(0).
 Although the main theorem and its application to asynchronous Q-learning are correct, TD(0) does not satisfy a key assumption for the main theorem [^3] In Chandak et al.
@@
 and hence the theorem was incorrectly applied to TD(0).
-We remove the need for that assumption in this version, giving a completely different proof specific to the TD(0) algorithm.
+We remove the need for that assumption in this version, giving a completely different proof tailored to the TD(0) algorithm.
 The previous paper required the iterates of the stochastic approximation iteration to be bounded by a constant with probability 1.
@@
 2023).
-The key difficulty caused by the lack of this assumption is in applying martingale inequalities which often require some restrictions on the increments of the martingale that are often not easy to verify.
-We do not work with the projection operator, and instead adapt relaxed concentration inequalities (Chung and Lu 2006, Section 8) for our problem.
+The key difficulty caused by the lack of this assumption is in applying martingale inequalities, which often require some restrictions on the increments of the martingale that are often not easy to verify.
+We do not modify the algorithm and instead adapt relaxed concentration inequalities (Chung and Lu 2006, Section 8) for our problem.
 These bounds have an extra term given by the probability of increments going above a certain threshold (Tao and Vu 2015, Prop.
 34).
-While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$.
-In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$.
-Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al.
+While the proof in this paper is restricted to TD(0), the underlying idea of using relaxed concentration inequalities is broadly applicable to other algorithms that face similar challenges due to unboundedness.
+1.1 Related Works
+There has been growing interest in analyzing the finite-time performance of reinforcement learning (RL) algorithms.
+Existing results can broadly be categorized by the type of bounds they establish.
+The most extensive body of work concerns expectation or mean-square bounds (see, e.g., Chen et al.
+(2020, 2021)).
+Another prominent line of research focuses on regret bounds, which characterize how the cumulative error grows over time—typically through almost-sure or expected regret guarantees (see, e.g., Azar et al.
+(2017), Jin et al.
+(2018), Yang et al.
+(2020), Yang and Wang (2019)).
+A third class comprises high-probability or concentration bounds (see, e.g., Li et al.
+(2023), Qu and Wierman (2020), Even-Dar and Mansour (2004).
+Our work falls within this category but differs from conventional analyses that establish high-probability guarantees only for sufficiently large time $n$.
+In contrast, we derive uniform all-time bounds, i.e., bounds that hold for all $n\geq n_{0}$ with probability at least $1-\delta$.
+Specifically for TD(0), moment bounds have been established in Srikant and Ying (2019), Bhandari et al.
 (2018), and Chen et al.
 (2021).
-Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al.
-(2021) (consider uniform sampling from dataset), Patil et al.
-(2023) (consider projection and tail averaging), and Chen et al.
-(2025), Dalal et al.
-(2018a) (consider i.i.d.
-samples of state - action - next state triplets $(s,a,s^{\prime})$).
-Going beyond TD(0), Bhandari et al.
-(2018), Chen et al.
-(2020, 2021), Wang et al.
-(2020), Dalal et al.
-(2018a, b), Li et al.
-(2020, 2023), Even-Dar and Mansour (2004), Prashanth et al.
-(2021), Qu and Wierman (2020), Sidford et al.
-(2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms.
-There is also a parallel body of work which seeks bound, either finite time or asymptotic, on the difference between the value function under the learned policy and the optimal value function (Jin et al.
-2018, Yang et al.
-2020, Yang and Wang 2019), i.e., the ‘regret’.
-Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
-Inspired by Chen et al.
-(2020), one of us considered stochastic approximation involving contractive maps and martingale noise, and derived such concentration bounds ‘from some $n_{0}$ on’ for this class of algorithms (Borkar 2021).
-This covered in particular synchronous Q-learning for discounted cost and some related schemes.
+High probability bounds have been established under various modifications of the TD(0) algorithm.
+These include uniform sampling from dataset (Prashanth et al.
+2021), projection and tail averaging (Patil et al.
+2023), and oracle access to i.i.d.
+samples of state - action - next state triplets $(s,a,s^{\prime})$ (Chen et al.
+2025, Dalal et al.
+2018).
+One of us considered stochastic approximation involving contractive maps and martingale noise, and derived maximal concentration bounds for this class of algorithms (Borkar 2021).
+This covered, in particular, synchronous Q-learning for discounted cost and some related schemes.
 In Chandak et al.
@@
 (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes.
-Another work which considers all-time bounds is Chen et al.
-(2025), but they only consider additive and multiplicative noise and not Markovian noise as considered in this paper.
-To summarize, conventional concentration or sample complexity bounds yield bounds, e.g.
-mean square or ‘with high probability’ bounds, on the error from the target after $n$ iterations starting from time zero.
-The asymptotic regret bounds get an upper bound or lower bound or both on how some measure of cumulative error grows with time in an asymptotic sense.
-Our bounds differ from both.
-They are ‘all time bounds’ in the sense that they give a high probability bound for the iterates to remain in a prescribed small neighborhood of the target for all time from some time $n_{0}$ onwards.
-In fact the requirement ‘from some $n_{0}$ onwards’ is dictated by the fact that the decreasing stepsize needs to be sufficiently small from $n_{0}$ on.
-As for regret bounds, these are for cumulative error and are typically ‘almost sure’ or ‘in mean’ and asymptotic in nature unlike our bounds, which are from some time on, but with high probability.
+A recent work (Chen et al.
+2025) considers all-time bounds for iterates without almost-sure boundedness, but they only consider additive and multiplicative noise and not Markovian noise as considered in this paper.
+Their proof technique relies on Moreau envelopes and a bootstrapping technique, which differs significantly from our work.
+1.2 Outline and Notation
 The rest of the paper is structured as follows.
-Section 2 gives a background to the TD(0) algorithm, and states the main result.
+Section 2 gives a background to the TD(0) algorithm, along with the required assumptions and the stochastic approximation formulation.
+Section 3 states the main result and provides some insights into the result.
 The result is proved in Section 4.
@@
 Appendix A states a martingale inequality used in our proof and Appendix B gives proofs for some technical lemmas which are used to prove the main theorem.
-1.1 Notation
 […]
-2 TD(0) - Background and Main Result
-TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies (Tsitsiklis and Van Roy 1997).
-Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $S$.
+2 Background on TD(0)
+TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies.
+Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $\mathcal{S}$.
 The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed.
-Assume that the chain is aperiodic irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$.
-Let $D$ denote the $s\times s$ diagonal matrix whose $i$th diagonal entry is $\pi(i)$.
-Reward $r(i)$ is received when a transition from state $i$ takes place.
+Assume that the chain is aperiodic irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(S)],S=|\mathcal{S}|$.
+Let $D$ denote the $S\times S$ diagonal matrix whose $s$th diagonal entry is $\pi(s)$.
+Reward $r(s)$ is received when a transition from state $s$ takes place.
+Note that this reward can be stochastic as well, and the additional noise thereof can be combined with other noise terms without affecting our concentration result.
+For simplicity, we assume that we receive a deterministic $r(s)$.
 The objective is to evaluate the long term discounted reward for each state given by the “value function”
-$V(i)=E\left[\sum_{m=0}^{\infty}\gamma^{m}r(X_{m})|X_{0}=i\right],i\in S.$
-Here $10<\gamma<$ 1 is a discount factor. The dynamic programming equation for evaluating the same is given by
-$V(i)=r(i)+\gamma\sum_{j}p(j|i)V(j),\ i\in S,$
-which can be written as the following vector equation
-$V=r+\gamma PV$
-for $r=[r(1),...,r(s)]^{T}$ and $P=[[p(j|i)]]_{i,j\in S}\in\mathcal{R}^{s\times s}$.
-The state space can often be large ($s\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}:S\mapsto\mathcal{R},1\leq i\leq d$, with $s>>d\geq 1$.
-Also, let $\varphi(i)=[\phi_{1}(i),\ldots,\phi_{d}(i)]^{T}$.
-Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\varphi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$.
+$V(s)=E\left[\sum_{m=0}^{\infty}\gamma^{m}r(X_{m})|X_{0}=s\right],s\in\mathcal{S}.$
+Here $0<\gamma<1$ is the discount factor. The dynamic programming equation for evaluating the same is
+$V(s)=r(s)+\gamma\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)V(s^{\prime}),\ s\in\mathcal{S}.$
+This can be written as the following vector equation
+$V=r+\gamma PV,$
+for $r=[r(1),...,r(S)]^{T}$ and $P=[[p(s^{\prime}|s)]]_{s^{\prime},s\in\mathcal{S}}\in\mathbb{R}^{S\times S}$.
+The state space can often be large ($S\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}\in\mathbb{R}^{S},1\leq i\leq d$, with $S>>d\geq 1$.
+Also, let $\varphi(s)=[\phi_{1}(s),\ldots,\phi_{d}(s)]^{T}$ for $s\in\mathcal{S}$ denote the vector comprising of components corresponding to state $s$ in each feature.
+Thus $V(s)\approx\sum_{i=1}^{d}x(i)\phi_{i}(s)$, i.e., $V(s)\approx x^{T}\varphi(s)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $S\times d$ matrix whose $i$th column is $\phi_{i}$.
 Here $x$ denotes the learnable weights for the linear function approximator.
@@
 where $\Pi$ denotes the projection to Range($\Phi$) with respect to a suitable norm.
-It turns out to be convenient to take projection with respect to the weighted norm $\|x\|_{D}\coloneqq\sqrt{x^{T}Dx}=(\sum_{i}\pi(i)|x(i)|^{2})^{1/2}$ whereby the projection map is
-$\Pi x\coloneqq\Phi(\Phi^{T}D\Phi)^{-1}\Phi^{T}Dx.$
+It turns out to be convenient to take projection with respect to the weighted norm $\|y\|_{D}\coloneqq\sqrt{y^{T}Dy}=(\sum_{s\in\mathcal{S}}\pi(s)(y(s))^{2})^{1/2}$ for $y\in\mathbb{R}^{S}$.
+The projection map with respect to this norm is
+$\Pi y\coloneqq\Phi(\Phi^{T}D\Phi)^{-1}\Phi^{T}Dy.$
 The invertibility of $\Phi^{T}D\Phi$ is guaranteed by the fact that $\Phi$ is full rank. Finally, the TD(0) algorithm is given by the recursion
 $x_{n+1}=x_{n}+a(n)\varphi(Y_{n})\Big(r(Y_{n})+\gamma\varphi(Y_{n+1})^{T}x_{n}-\varphi(Y_{n})^{T}x_{n}\Big).$ (2)
-Recall that $\varphi(i)^{T}\in\mathcal{R}^{d}$ for $i\in S$ denotes the $i$th row of $\Phi$ and $a(n)$ denotes the positive stepsize sequence. We now make the following two assumptions:
+Here $a(n)$ denotes the positive stepsize sequence. At the end of this section, we explain how this iteration can be expected to converge to the required fixed point from (1).
+2.1 Assumptions
+We impose two assumptions on the algorithm.
+The first is about the feature vectors, and as we explain next, it does not restrict the algorithm.
+The second specifies the class of stepsizes $a(n)$ considered, which is standard in analysis of RL.
+In fact, our results hold for a broader class of step sizes than those typically required in stochastic approximation frameworks.
   1. 1.
-    First we make an assumption on $\Phi$.
-For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the square of the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$.
+    For the assumption on $\Phi$, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the square of the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$.
 Assume that
@@
 This assumption can be replaced with the following assumption on the basis vectors.
-    $\|\varphi(i)\|\leq\frac{\sqrt{2(1-\gamma)}}{1+\gamma}\;\forall i,$
-    i.e., the $\ell_{2}$ norm of each row of $\Phi$ is bounded by $\sqrt{2(1-\gamma)}/1+\gamma$. To see this, note that
-    $\frac{\|\Psi x\|_{2}}{\|x\|_{2}}=\frac{\|\Phi x\|_{D}}{\|x\|_{2}}\leq\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}.$
-    Now $\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}$ is the operator norm defined with $\ell_{2}$ norm for domain and $\ell_{\infty}$ for co-domain, which is equal to the maximum $\ell_{2}$ norm of a row.
-Hence
-    $\lambda_{M}=\max_{x\neq\theta}\frac{\|\Psi x\|_{2}}{\|x\|_{2}}\leq\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}=\max_{i}\|\varphi(i)\|_{2}.$
+∥φ(s)∥≤2(1-γ)1+γ ∀s∈S, i.e., the $\ell_{2}$ norm of each row of $\Phi$ is bounded by $\sqrt{2(1-\gamma)}/(1+\gamma)$.
+To see this, note that ∥Ψx∥2∥x∥2=∥Φx∥D∥x∥2≤∥Φx∥∥x∥2.Now $\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}$ is the operator norm defined with $\ell_{2}$ norm for domain and $\ell_{\infty}$ for co-domain, which is equal to the maximum $\ell_{2}$ norm of a row.
+Hence λ_M=max_x≠θ∥Ψx∥2∥x∥2≤max_x≠θ∥Φx∥∥x∥2=max_s∈Sφ(s)∥_2.
   2. 2.
@@
 The above assumptions on the stepsize sequence can be weakened so as to hold only after some $N>1$, without any changes in the analysis.
-We next rearrange iteration (2) to separate the martingale noise and the ‘Markov noise’.
+2.2 Formulation as a Stochastic Approximation Iteration
+We next rearrange algorithm (2) to separate the martingale noise and the ‘Markov noise’, and write it as a stochastic approximation iteration.
 $x_{n+1}$ $=$ $x_{n}+a(n)\varphi(Y_{n})\Big(r(Y_{n})+\gamma\varphi(Y_{n+1})^{T}x_{n}-\varphi(Y_{n})^{T}x_{n}\Big)$
 $=$ $x_{n}+a(n)\bigg(F(x_{n},Y_{n})-x_{n}+M_{n+1}x_{n}\bigg)$
-$=$ $x_{n}+a(n)\left(\sum_{i\in S}\pi(i)F(x_{n},i)-x_{n}\right)+\underbrace{a(n)M_{n+1}x_{n}}_{\tau_{1}}+\underbrace{a(n)\left(F(x_{n},Y_{n})-\sum_{i\in S}\pi(i)F(x_{n},i)\right)}_{\tau_{2}}.$
+$=$ $x_{n}+a(n)\left(\sum_{s\in\mathcal{S}}\pi(s)F(x_{n},s)-x_{n}\right)+\underbrace{a(n)M_{n+1}x_{n}}_{\tau_{1}}+\underbrace{a(n)\left(F(x_{n},Y_{n})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{n},s)\right)}_{\tau_{2}},$
 where
-$F(x,Y)=\varphi(Y)r(Y)+\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}x-\varphi(Y)\varphi(Y)^{T}x+x$
+$F(x,Y)=\varphi(Y)r(Y)+\gamma\varphi(Y)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y)\varphi(s^{\prime})^{T}x-\varphi(Y)\varphi(Y)^{T}x+x,$
 and
-$M_{n+1}=\gamma\varphi(Y_{n})\Big{(}\varphi(Y_{n+1})^{T}-\sum_{j}p(j|Y_{n})\varphi(j)^{T}\Big{)}.$
+$M_{n+1}=\gamma\varphi(Y_{n})\Big(\varphi(Y_{n+1})^{T}-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{n})\varphi(s^{\prime})^{T}\Big).$
 Define the family of $\sigma$-fields $\mathcal{F}_{n}\coloneqq\sigma(x_{0},Y_{m},m\leq n),\ n\geq 0$.
@@
 $E[M_{n+1}x_{n}|\mathcal{F}_{n}]=\theta,a.s.\;\forall n,$
-where $\theta$ denotes the zero vector. Then term $\tau_{1}$ denotes the error due to the martingale noise term and term $\tau_{2}$ denotes the error due to the ‘Markov noise’ $\{Y_{n}\}$.
-The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction.
+where $\theta$ denotes the zero vector. The term $\tau_{1}$ denotes the error due to the martingale noise term and term $\tau_{2}$ denotes the error due to the ‘Markov noise’ $\{Y_{n}\}$.
+The following lemma shows that the function $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$ is a contraction.
 Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$.
-Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor.
+While the contraction property of the TD(0) algorithm is well known Tsitsiklis and Van Roy (1997), we obtain an explicit expression for the contraction factor.
 Lemma 1.
-For any $x,z\in\mathcal{R}^{d}$,
-$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|\leq\alpha\|x-z\|,$
+For any $x,z\in\mathbb{R}^{d}$,
+$\|\sum_{s\in\mathcal{S}}\pi(s)(F(x,s)-F(z,s))\|\leq\alpha\|x-z\|,$
 where
 $\alpha=\sqrt{1-\min_{x\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg(2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg)}.$
-Moreover, $10<\alpha<$ 1 and hence the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction.
+Moreover, $0<\alpha<1$ and hence the function $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$ is a contraction.
 The proof appears in Appendix B.
-The Banach contraction mapping theorem implies that $\sum_{i}\pi(i)F(\cdot,i)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{i}\pi(i)F(x^{*},i)=x^{*}$.
+The Banach contraction mapping theorem implies that $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{s\in\mathcal{S}}\pi(s)F(x^{*},s)=x^{*}$.
+We next show that the fixed point $x^{*}$ is the required fixed point we wish to converge to.
+Before that, we first observe that
+$\sum_{s\in\mathcal{S}}\pi(s)F(x,s)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x.$
 Then,
-$\sum_{i}\pi(i)F(x^{*},i)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x^{*}=x^{*}$
+$\sum_{s\in\mathcal{S}}\pi(s)F(x^{*},s)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x^{*}=x^{*}$
 $\implies(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi)x^{*}=\Phi^{T}D\Phi x^{*}$
@@
 So $\Phi x^{*}$ is the required fixed point of (1).
-For $n\geq 0$, we further define:
-[…]
+3 Main Result
+Before stating the main result, we define the following two sequences. For $n\geq 0$,
+$b_{k}(n)$ $=$ $\sum_{m=k}^{n}a(m),\ 0\leq k\leq n<\infty,$
 $\beta_{k}(n)$ $=$ $\begin{cases}\frac{1}{k^{d_{2}-d_{1}}n^{d_{1}}},&\text{if}\ d_{1}\leq d_{2}\\ \frac{1}{n^{d_{2}}},&\text{otherwise}.\end{cases}$
@@
 Theorem 1.
-There exist finite positive constants $c_{1},c_{2},c_{3}$ and $D$ such that for $0<\delta\leq 1,0<\epsilon\leq 1$, $n_{0}>0$ large enough to satisfy $\alpha+a(n_{0})c_{1}<1$, and $n\geq n_{0}$,
+There exist finite positive constants $c_{1},c_{2}$ and $D$ such that for $0<\delta\leq 1,0<\epsilon\leq 1$, $n_{0}>0$ large enough to satisfy $\alpha+a(n_{0})c_{1}<1$, and $n\geq n_{0}$,
   1. (a)
-    the inequality
-    $\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n,$
-    holds with probability exceeding
-    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
+    the inequality x_m-x^*∥≤e^-(1-α)b_n_0(m-1)ϵ+a(n0)(c2+c1ϵ)+δ1-α-a(n0)c1,   n_0mn, holds with probability exceeding 1-2d_m=n_0+1^n e^-Dδ^2/β_n_0(m-1)-P(x_n_0-x^*¿ϵ).
   2. (b)
-    In particular,
-    $\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall m\geq n_{0},$
-    holds with probability exceeding
-    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
+    In particular, x_m-x^*∥≤e^-(1-α)b_n_0(m-1)ϵ+a(n0)(c2+c1ϵ)+δ1-α-a(n0)c1,   mn_0, holds with probability exceeding 1-2d_mn_0+1 e^-Dδ^2/β_n_0(m-1n)-P(x_n_0-x^*¿ϵ).
 The following are some remarks about the theorem and the proof that follows:
@@
 Define $z_{n}$ for $n\geq n_{0}$ by
-$z_{n+1}=z_{n}+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-z_{n}\right),$
+$z_{n+1}=z_{n}+a(n)\left(\sum_{s}\pi(s)F(z_{n},s)-z_{n}\right),$
 where $z_{n_{0}}=x_{n_{0}}$. Note that $\|x_{n}-x^{*}\|\leq\|x_{n}-z_{n}\|+\|z_{n}-x^{*}\|$. To bound the second term, note that
-$z_{n+1}-x^{*}$ $=(1-a(n))(z_{n}-x^{*})+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-x^{*}\right)$
-$=(1-a(n))(z_{n}-x^{*})+a(n)\sum_{i}\pi(i)\left(F(z_{n},i)-F(x^{*},i)\right).$
-The second equality follows from the fact that $x^{*}$ is a fixed point for $\sum_{i}\pi(i)F(\cdot,i)$. Then,
-$\|z_{n+1}-x^{*}\|$ $\leq(1-a(n))\|z_{n}-x^{*}\|+a(n)\|\sum_{i\in S}\pi_{x_{n}}(i)(F(z_{n},i)-F(x^{*},i))\|$
+$z_{n+1}-x^{*}$ $=(1-a(n))(z_{n}-x^{*})+a(n)\left(\sum_{s\in\mathcal{S}}\pi(s)F(z_{n},s)-x^{*}\right)$
+$=(1-a(n))(z_{n}-x^{*})+a(n)\sum_{s\in\mathcal{S}}\pi(s)\left(F(z_{n},s)-F(x^{*},s)\right).$
+The second equality follows from the fact that $x^{*}$ is a fixed point for $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$. Then,
+$\|z_{n+1}-x^{*}\|$ $\leq(1-a(n))\|z_{n}-x^{*}\|+a(n)\|\sum_{s\in\mathcal{S}}\pi(s)(F(z_{n},s)-F(x^{*},s))\|$
 $\leq(1-(1-\alpha)a(n))\|z_{n}-x^{*}\|.$
 which finally gives us
-$\|z_{n}-x^{*}\|\leq\psi(n,n_{0})\|x_{n_{0}}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(n-1)}\|x_{n_{0}}-x^{*}\|,$ (4)
-where $\psi(n,m)\coloneqq\prod_{k=m}^{n-1}(1-(1-\alpha)a(k))$ if $n>m$ and $1$ otherwise. This also implies that for all $n\geq n_{0}$
+$\|z_{n}-x^{*}\|\leq\prod_{k=n_{0}}^{n-1}(1-(1-\alpha)a(k))\|x_{n_{0}}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(n-1)}\|x_{n_{0}}-x^{*}\|,$ (4)
+This also implies that for all $n\geq n_{0}$,
 $\|z_{n}\|\leq\|x_{n_{0}}-x^{*}\|+\|x^{*}\|.$ (5)
@@
 $x_{n+1}-z_{n+1}=$ $(1-a(n))(x_{n}-z_{n})+a(n)M_{n+1}x_{n}$
-$+a(n)\left(F(x_{n},Y_{n})-\sum_{i}\pi(i)F(z_{n},i)\right)$
+$+a(n)\left(F(x_{n},Y_{n})-\sum_{s}\pi(s)F(z_{n},s)\right)$
 $=(1-a(n))(x_{n}-z_{n})+\ a(n)M_{n+1}x_{n}$
-$+\ a(n)\left(\sum_{i}\pi(i)(F(x_{n},i)-F(z_{n},i))\right)$
-$+\ a(n)\left(F(x_{n},Y_{n})-\sum_{i}\pi(i)F(x_{n},i)\right).$
+$+\ a(n)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{n},s)-F(z_{n},s))\right)$
+$+\ a(n)\left(F(x_{n},Y_{n})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{n},s)\right).$
 For $n,m\geq 0$, let $\chi(n,m)=\prod_{k=m}^{n}(1-a(k))$ if $n\geq m$ and $1$ otherwise. For some $n\geq n_{0}$, we iterate the above for $n_{0}\leq m\leq n$ to obtain
 $x_{m+1}-z_{m+1}=$ $\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)M_{k+1}x_{k}$
-$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)$
-$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right).$ (6)
+$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{k},s)-F(z_{k},s))\right)$
+$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right).$ (6)
 Here we use the definition that $x_{n_{0}}=z_{n_{0}}$. We first simplify the third term above. For simplicity, we define $F(x,Y)=F_{1}(Y)+F_{2}(Y)x+x,$ where
-$F_{1}(Y)=\varphi(Y)r(Y)\in\mathcal{R}^{d}\;\;\;\textrm{and}\;\;\;F_{2}(Y)=\left(\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}-\varphi(Y)\varphi(Y)^{T}\right)\in\mathcal{R}^{d\times d}.$
-We define $U:S\mapsto\mathcal{R}^{d}$ to be a solution of the Poisson equation:
-$U(i)=F_{1}(i)-\sum_{j}\pi(j)F_{1}(j)+\sum_{j}p(j|i)U(j),\ i\in S.$ (7)
-For $i_{0}\in S$, $\tau\coloneqq\min\{n>0:Y_{n}=i_{0}\}$ and $E_{i}[\cdots]=E[\cdots|Y_{0}=i]$, we know that
-$U^{\prime}(i)=E_{i}\left[\sum_{m=0}^{\tau-1}(F_{1}(Y_{m})-\sum_{j\in\mathcal{S}}\pi(j)F_{1}(j))\right],\ i\in S,$ (8)
+$F_{1}(Y)=\varphi(Y)r(Y)\in\mathbb{R}^{d}\;\;\;\textrm{and}\;\;\;F_{2}(Y)=\left(\gamma\varphi(Y)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y)\varphi(s^{\prime})^{T}-\varphi(Y)\varphi(Y)^{T}\right)\in\mathbb{R}^{d\times d}.$
+We define $U:\mathcal{S}\mapsto\mathbb{R}^{d}$ to be a solution of the Poisson equation:
+$U(s)=F_{1}(s)-\sum_{s^{\prime}\in\mathcal{S}}\pi(s^{\prime})F_{1}(s^{\prime})+\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)U(s^{\prime}),\ s\in\mathcal{S}.$ (7)
+For $s_{0}\in\mathcal{S}$, $\tau\coloneqq\min\{n>0:Y_{n}=s_{0}\}$ and $E_{s}[\cdots]=E[\cdots|Y_{0}=s]$, we know that
+$U^{\prime}(s)=E_{s}\left[\sum_{m=0}^{\tau-1}(F_{1}(Y_{m})-\sum_{s^{\prime}\in\mathcal{S}}\pi(s^{\prime})F_{1}(s^{\prime}))\right],\ s\in S,$ (8)
 is one particular solution to the Poisson equation (see, e.g., Lemma 4.2 and Theorem 4.2 of Section VI.4, pp.
 85-91, of Borkar (1991)).
-Thus $\|U^{\prime}(i)\|_{\infty}\leq 2\max_{i}\|F_{1}(i)\|_{\infty}E_{i}[\tau]$.
-For an irreducible Markov chain with a finite state space, $E_{i}[\tau]$ is finite for all $i$ and hence the solution $U^{\prime}(i)$ is bounded for all $i$.
+Thus $\|U^{\prime}(s)\|_{\infty}\leq 2\max_{s\in\mathcal{S}}\|F_{1}(s)\|_{\infty}E_{s}[\tau]$.
+For an irreducible Markov chain with a finite state space, $E_{s}[\tau]$ is finite for all $s$ and hence the solution $U^{\prime}(s)$ is bounded for all $s$.
 For each $\ell$, the Poisson equation specifies $U^{\ell}(\cdot)$ uniquely only up to an additive constant.
-Along with the additional constraint that $U(i_{0})=0$ for a prescribed $i_{0}\in S$, the system of equations given by (7) has a unique solution.
-Henceforth $U$ refers to the unique solution of the Poisson equation with $U(i_{0})=0$.
-Similarly, let $W:S\mapsto\mathcal{R}^{d\times d}$ be a solution of the Poisson equation:
-$W(i)=F_{2}(i)-\sum_{j}\pi(j)F_{2}(j)+\sum_{j}p(j|i)W(j),\ i\in S,$ (9)
-with the additional constraint that $W(i_{0})=0$ for a prescribed $i_{0}\in S$ as above.
+Along with the additional constraint that $U(s_{0})=0$ for a prescribed $s_{0}\in S$, the system of equations given by (7) has a unique solution.
+Henceforth $U$ refers to the unique solution of the Poisson equation with $U(s_{0})=0$.
+Similarly, let $W:\mathcal{S}\mapsto\mathbb{R}^{d\times d}$ be the unique solution of the Poisson equation:
+$W(s)=F_{2}(s)-\sum_{s^{\prime}}\pi(s^{\prime})F_{2}(s^{\prime})+\sum_{s^{\prime}}p(s^{\prime}|s)W(s),\ s\in S,$ (9)
+with the additional constraint that $W(s_{0})=0$ for a prescribed $s_{0}\in S$ as above.
 The following lemma gives a simplification of the third term in (4), using the solutions of Poisson equation stated above.
@@
 There exist positive constants $c_{1},c_{2}$ such that for all $n_{0}\leq m\leq n$,
-$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)$
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right)$
 $=\ \sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\widetilde{U}_{k+1}+\widetilde{W}_{k+1}x_{k}\right)+\mu_{m}(n_{0}),$
@@
 $\|\mu_{m}(n_{0})\|\leq a(n_{0})(c_{2}+c_{1}x_{m}^{\prime}+c_{1}\|x_{n_{0}}-x^{*}\|).$
-Here $\widetilde{U}_{k+1}$ and $\widetilde{W}_{k+1}x_{k}$ are martingale difference sequences with respect to $\mathcal{F}_{k}$ where $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)$ and $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{j}p(j|Y_{k})W(j)$ for $k\geq n_{0}$ and and the zero vector, resp.
+Here $\widetilde{U}_{k+1}$ and $\widetilde{W}_{k+1}x_{k}$ are martingale difference sequences with respect to $\mathcal{F}_{k}$ where $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{s^{\prime}}p(s^{\prime}|Y_{k})U(s^{\prime})$ and $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{s^{\prime}}p(s^{\prime}|Y_{k})W(s^{\prime})$ for $k\geq n_{0}$ and the zero vector, resp.
 the zero matrix, otherwise.
 The proof appears in Appendix B. Returning to (4), we now have
-$x_{m+1}-z_{m+1}$ $=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)$
-$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)+\mu_{m}(n_{0})$
+$x_{m+1}-z_{m+1}$ $=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{k},s)-F(z_{k},s))\right)$
+$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)+\mu_{m}(n_{0}).$
 Now,
-$\|x_{m+1}-z_{m+1}\|$ $\leq\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)\right\|$
+$\|x_{m+1}-z_{m+1}\|$ $\leq\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{k},s)-F(z_{k},s))\right)\right\|$
 $\;\;\;+\ \left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|$
@@
 This implies that
-[…]
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\ \leq\ 1.$
 Using the definition of $x^{\prime}_{m}$, we have
 $x^{\prime}_{m+1}\leq$ $\;(\alpha+a(n_{0})c_{1})x^{\prime}_{m}+\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|$
-$+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
-Next we obtain a bound on the probability
-$P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right),$
-for some $\epsilon>0$ and $\delta>0$ (recall the assumption that $\alpha+a(n_{0})c_{1}<1$).
+$+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$ (11)
+Next we wish to obtain a bound on the probability
+$P\Big(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\Delta(n_{0},\epsilon,\delta),\;\forall n_{0}\leq m\leq n\Big),$
+for some $\epsilon>0$ and $\delta>0$ (recall the assumption that $\alpha+a(n_{0})c_{1}<1$). For ease of notation, here we have defined
+$\Delta(n_{0},\epsilon,\delta)\coloneqq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
 From (4), recall that $\|z_{n}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(n-1))\|x_{n_{0}}-x^{*}\|\;a.s.,$ and hence,
-$\|x_{n_{0}}-x^{*}\|\leq\epsilon\implies\|z_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|.$
+$\|x_{n_{0}}-x^{*}\|\leq\epsilon\implies\|z_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon.$
 Also recall that $\sup_{n_{0}\leq m\leq n}\|x_{m}-z_{m}\|=x_{n}^{\prime}$. Hence
-$\left\{\|x_{n_{0}}-x^{*}\|\leq\epsilon\right\}\bigcap\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}$
-$\subseteq\left\{\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right\}.$
+$\left\{\|x_{n_{0}}-x^{*}\|\leq\epsilon\right\}\bigcap\left\{x^{\prime}_{n}\leq\Delta(n_{0},\epsilon,\delta)\right\}$
+$\subseteq\left\{\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\Delta(n_{0},\epsilon,\delta),\;\forall n_{0}\leq m\leq n\right\}.$
 This implies the following relation between the probabilities of the two sets.
-$P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right)$
-$\geq 1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
+$P\Big(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|+\Delta(n_{0},\epsilon,\delta),\;\forall n_{0}\leq m\leq n\Big)$
+$\geq 1-P\left(\Big\{x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\Big\}\bigcup\Big\{\|x_{n_{0}}-x^{*}\|>\epsilon\Big\}\right).$
 To compensate for the lack of an almost sure bound on the iterates $\{x_{n}\}$, we adapt the proof method from Proposition 34 from Tao and Vu (2015) (see Section 8 of Chung and Lu (2006) for a detailed explanation).
 For this, we define $\xi=\{x_{0},Y_{k},k\geq 0\}$ and the ‘bad’ set $\mathbb{B}_{m}$ as
-$\mathbb{B}_{m}=\left\{\xi\mid x_{m}^{\prime}(\xi)>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}(\xi)-x^{*}\|>\epsilon\right\}.$
+$\mathbb{B}_{m}=\left\{\xi\mid x_{m}^{\prime}(\xi)>\Delta(n_{0},\epsilon,\delta)\;\;\bigcup\;\;\|x_{n_{0}}(\xi)-x^{*}\|>\epsilon\right\}.$
 Here the notation $x_{m}^{\prime}(\xi)$ and $x_{n_{0}}(\xi)$ highlights the dependence of $x_{m}^{\prime}$ and $x_{n_{0}}$ on the realizations of $x_{0}$ and $\{Y_{k}\}$.
@@
 Also, define $\bar{x}_{m,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq m}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$.
-Now, adapting the proof of Proposition 34 from Tao and Vu (2015) for our case, we first note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies
-$P(\bar{x}_{n,n-1}^{\prime}(\xi)\neq x_{n}^{\prime}(\xi))\leq P(\mathbb{B}_{n-1}).$
-Henceforth we drop $\xi$ for ease of notation, rendering implicit the dependence of all random variables on $\xi$. Then,
-$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
-$\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\stackrel{{\scriptstyle(a)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\stackrel{{\scriptstyle(b)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\stackrel{{\scriptstyle(c)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
-$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
+Note that, $\bar{x}^{\prime}_{m,n-1}=0$ when $\xi\in\mathbb{B}_{n-1}$, and $\bar{x}^{\prime}_{m,n-1}=x^{\prime}_{m}\leq\Delta(n_{0},\epsilon,\delta)$ when $\xi\notin\mathbb{B}_{n-1}$.
+The intuition behind these definitions is that $\bar{x}^{\prime}_{m,n-1}$ is always bounded by $\Delta(n_{0},\epsilon,\delta)$ for all $m\leq n-1$.
+Note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies $P(\bar{x}_{n,n-1}^{\prime}(\xi)\neq x_{n}^{\prime}(\xi))\leq P(\mathbb{B}_{n-1}).$ Henceforth we drop $\xi$ for ease of notation, rendering implicit the dependence of all random variables on $\xi$.
+Then,
+$P\left(x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\right)$ $\leq P\left(x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\stackrel{{\scriptstyle(c)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)+P(\mathbb{B}_{n-1})$
+$=P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)+P\left(x_{n-1}^{\prime}>\Delta(n_{0},\epsilon,\delta)\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
 Inequality (a) here follows from the observation that
-$\left\{\bar{x}^{\prime}_{n,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcap\{\bar{x}^{\prime}_{n,n-1}=x^{\prime}_{n}\}\subseteq\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\},$
+$\left\{\bar{x}^{\prime}_{n,n-1}\leq\Delta(n_{0},\epsilon,\delta)\right\}\bigcap\{\bar{x}^{\prime}_{n,n-1}=x^{\prime}_{n}\}\subseteq\left\{x^{\prime}_{n}\leq\Delta(n_{0},\epsilon,\delta)\right\},$
 which implies that
-$\left\{x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\subseteq\left\{\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcup\{\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\},$
+$\left\{x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\right\}\subseteq\left\{\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right\}\bigcup\{\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\},$
 which gives us the required inequality.
 Inequality (b) follows from union bound and inequality (c) follows from the observations that $\{\|x_{n_{0}}-x^{*}\|>\epsilon\}\subseteq\mathbb{B}_{n-1}$ and $\{\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\}\subseteq\mathbb{B}_{n-1}$.
-Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction.
-Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$.
-Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$.
-Then,
-$P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)=P\left(\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right).$
-We can use (3) to obtain a bound on this probability.
-In the case where $\xi_{n-1}\in\mathbb{B}_{n-1}$, $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|=0$ and in the case where $\xi_{n-1}\notin\mathbb{B}_{n-1}$, we have $x_{k}=\bar{x}_{k,n-1}$.
-So (3) can be written as:
+Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)$ by induction.
+We first note that $\bar{x}^{\prime}_{n-1,n-1}$ is bounded by $\Delta(n_{0},\epsilon,\delta)$ by definition.
+Hence $P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)=P\left(\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|>\Delta(n_{0},\epsilon,\delta)\right)$.
+We first restate (4) for m=n1m=n-1.
+$\|x_{n}-z_{n}\|\leq x^{\prime}_{n}\leq$ $\;(\alpha+a(n_{0})c_{1})x^{\prime}_{n-1}+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|)$
+$+\left\|\sum_{k=n_{0}}^{n-1}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|.$
+Now, let $I\{\cdot\}$ denote the indicator function which is $1$ when $\{\cdot\}$ holds true, and zero otherwise.
+$\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$
+$=\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|I\{\xi_{n-1}\in\mathbb{B}_{n-1}\}+\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}$
+$\stackrel{{\scriptstyle(a)}}{{=}}0\times I\{\xi_{n-1}\in\mathbb{B}_{n-1}\}+\|x_{n}-z_{n}\|\times I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}$
+$\leq I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}\times\Bigg((\alpha+a(n_{0})c_{1})x^{\prime}_{n-1}+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|)$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|\Bigg)$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}\times\Bigg((\alpha+a(n_{0})c_{1})\Delta(n_{0},\epsilon,\delta)+a(n_{0})(c_{2}+c_{1}\epsilon)$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|\Bigg).$
+Here inequality (a) follows from our definition of $\mathbb{B}_{n-1}$ that $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|=0$ when $\xi_{n-1}\in\mathbb{B}_{n-1}$, and $x_{n}=\bar{x}_{n,n-1}$ when $\xi_{n-1}\notin\mathbb{B}_{n-1}$.
+Inequality (b) follows from the fact that when $\xi_{n-1}\notin\mathbb{B}_{n-1}$, then $x_{k}=\bar{x}_{k,n-1}$ for all $k$, and $\|x_{n_{0}}-x^{*}\|\leq\epsilon$.
+Substituting the expression for $\Delta(n_{0},\epsilon,\delta)$ we obtain the following.
 $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$ $\leq(\alpha+a(n_{0})c_{1})\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}+a(n_{0})(c_{2}+c_{1}\epsilon)$
@@
 $\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)}{1-\alpha-a(n_{0})c_{1}}+\frac{\alpha+a(n_{0})c_{1}}{1-\alpha-a(n_{0})c_{1}}\delta$
-$\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|$
+$\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|.$
 When
@@
 $\leq P\left(\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|>\delta\right).$
-Let us denote the probability on the right side of the inequality as $p_{n}$. Then
-$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq p_{n}+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
+Let us denote the probability on the right side of the inequality as $p_{n-1}$. Then
+$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq p_{n-1}+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
 Then repeating the same procedure using $\mathbb{B}_{n-2}$, we obtain
 $P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
-$\leq p_{n-1}+P\left(x_{n-2}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
+$\leq p_{n-2}+P\left(x_{n-2}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
 Iterating this for $n\geq m\geq n_{0}+1$, we get
-$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq\sum_{m={n_{0}+1}}^{n}p_{m}+P(\|x_{n_{0}}-x^{*}\|>\epsilon).$
+$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq\sum_{m={n_{0}+1}}^{n}p_{m-1}+P(\|x_{n_{0}}-x^{*}\|>\epsilon).$
+The probabilities $p_{m}$ can be bounded using standard martingale inequalities as the terms of the martingale difference sequence are almost surely bounded.
 The following lemma, proved in Appendix B, gives a bound on the probabilities $p_{m}$:
 Lemma 3.
-There exist positive constants $c_{3}$ and $D$ such that for $0<\epsilon\leq 1,0<\delta\leq 1$,
-$p_{m}\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)},0<\delta\leq C,$
-$p_{m}\leq 2de^{-D\delta/\beta_{n_{0}}(m)},\delta>C,$
-where $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$. Recall that $d$ here denotes the dimension of the iterates $\{x_{n}\}$.
+There exists positive constant $D$ such that for $0<\epsilon\leq 1,0<\delta\leq 1$,
+$p_{m}\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)}.$
+Recall that $d$ here denotes the dimension of the iterates $\{x_{n}\}$.
 This completes the proof for the first part of Theorem 1.
@@
 Let $\{M_{n}\}$ be a real valued martingale difference sequence with respect to an increasing family of $\sigma$-fields $\{\mathcal{F}_{n}\}$. Assume that there exist $\varepsilon,C>0$ such that
-[…]
+$E\left[e^{\varepsilon|M_{n}|}\Big|\mathcal{F}_{n-1}\right]\leq C\ \ \forall\;n\geq 1,\mbox{a.s.}$
 Let $S_{n}:=\sum_{m=1}^{n}\zeta_{m,n}M_{m}$, where $\zeta_{m,n},\ m\leq n,$, for each $n$, are a.s.
@@
 $P\left(|S_{n}|>\epsilon\right)$ $\leq$ $2e^{-\frac{D\epsilon^{2}}{\omega(n)}},\ \ \mbox{if}\ \epsilon\in\left(0,\frac{C\gamma_{1}}{\varepsilon}\right],$ (13)
-$2e^{-\frac{D\epsilon}{\omega(n)}},\ \ \mbox{otherwise.}$.
+$2e^{-\frac{D\epsilon}{\omega(n)}},\ \ \mbox{otherwise.}$
 This is a variant of Theorem 1.1 of Liu and Watbled (2009). See Thoppe and Borkar (2019), Theorem A.1, pp. 21-23, for details.
@@
 Proof.
-$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$ $=$ $\|\gamma\sum_{i}\pi(i)\varphi(i)\sum_{j}p(j|i)\varphi(j)^{T}(x-z)$ (13)
-$\;\;\;\;\;\;\;-\sum_{i}\pi(i)\varphi(i)\varphi(i)^{T}(x-z)+(x-z)\|^{2}$
+$\|\sum_{s\in\mathcal{S}}\pi(s)(F(x,s)-F(z,i))\|^{2}$ $=$ $\|\gamma\sum_{s\in\mathcal{S}}\pi(s)\varphi(s)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)\varphi(s^{\prime})^{T}(x-z)$ (14)
+$\;\;\;\;\;\;\;-\sum_{s\in\mathcal{S}}\pi(s)\varphi(s)\varphi(s)^{T}(x-z)+(x-z)\|^{2}$
 $=$ $\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)(x-z)\|^{2}$
@@
 $=$ $2\|\Phi(x-z)\|_{D}^{2}.$
-Inequality (a) follows from the Cauchy-Schwarz inequality and (b) follows from the observation that $\|Px\|_{D}\leq\|x\|_{D}$, which can be proved as follows.
-$\|Px\|^{2}_{D}=\sum_{i\in S}\pi(i)\left(\sum_{j\in S}p(j|i)x_{j}\right)^{2}\leq\sum_{i\in S}\pi(i)\sum_{j\in S}p(j|i)x_{j}^{2}=\sum_{j\in S}\pi(j)x_{j}^{2}=\|x\|^{2}_{D}.$
+Inequality (a) follows from the Cauchy-Schwarz inequality and (b) follows from the observation that $\|Py\|_{D}\leq\|y\|_{D}$, which can be proved as follows.
+$\|Py\|^{2}_{D}=\sum_{s\in\mathcal{S}}\pi(s)\left(\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)y(s^{\prime})\right)^{2}\leq\sum_{s\in\mathcal{S}}\pi(s)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)y(s^{\prime})^{2}=\sum_{s^{\prime}\in\mathcal{S}}\pi(s^{\prime})y(s^{\prime})^{2}=\|y\|^{2}_{D}.$
 Here the inequality follows from Jensen’s inequality.
 Combining (15) and (16) with (14) gives us:
-$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$$\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}$ (16)
+$\|\sum_{s\in\mathcal{S}}\pi(s)(F(x,s)-F(z,s))\|^{2}$ $\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}$ (17)
 $+\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}.$
@@
 Thus
-$\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}$
-$=\|\Phi^{T}\sqrt{D}(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
+$\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}$ $=\|\Phi^{T}\sqrt{D}(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
 $\leq\lambda_{M}^{2}\|(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
@@
 The last inequality follows from the triangle inequality. We now invoke assumption (3) and combine (B.1) with (17) as follows:
-$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$$\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}+\lambda_{M}^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}$ (18)
+$\|\sum_{s\in\mathcal{S}}\pi(s)(F(x,s)-F(z,s))\|^{2}$ $\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}+\lambda_{M}^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}$ (19)
 $<$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}$
@@
 Note that as the columns of $\Phi$ are linearly independent, $x\neq\theta\implies\Phi x\neq\theta$ and hence $\frac{\|\Phi x\|_{D}}{\|x\|}>0$ when $x\neq\theta$.
-Also, note that $\min_{x\neq\theta}\frac{\|\Phi x\|_{D}}{\|x\|}=\min_{\|x\|=1}\|\Phi x\|_{D}$ and hence by extreme value theorem, we have that $\min_{\|x\|=1}\|\Phi x\|_{D}$ is attained and is greater than 00.
+Also, note that $\min_{x\neq\theta}\frac{\|\Phi x\|_{D}}{\|x\|}=\min_{\|x\|=1}\|\Phi x\|_{D}$ and hence by extreme value theorem, we have that $\min_{\|x\|=1}\|\Phi x\|_{D}$ is attained and is greater than 0.
 Along with assumption (3), this implies that $\alpha<1$.
@@
 Using definitions of $U(\cdot)$ and $W(\cdot)$ we have,
-$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)$
-$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{j}p(j|Y_{k})U(j)\right)$ (19a)
-$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{j}p(j|Y_{k})W(j)\right)x_{k}.$ (19b)
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right)$
+$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})\right)$ (20a)
+$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})W(s^{\prime})\right)x_{k}.$ (20b)
 We first simplify (20a) as follows:
-$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{j}p(j|Y_{k})U(j)\right)$
-$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)\right)$ (20a)
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})\right)$
+$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})\right)$ (21a)
 $\;\;\;+\sum_{k=n_{0}+1}^{m}\left((\chi(m,k+1)a(k)-\chi(m,k)a(k-1))U(Y_{k})\right)$ (21b)
 $\;\;\;+\chi(m,n_{0}+1)a(n_{0})U(Y_{n_{0}})-\chi(m,m+1)a(m)U(Y_{m+1}).$ (21c)
-For (20a), define $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)$ for $k\geq n_{0}$ and 00 otherwise. This is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
+For (21a), define $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})$ for $k\geq n_{0}$ and 0 otherwise.
+This is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
 We define $U_{max}\coloneqq\max_{i\in S}\|U(i)\|$ and bound the norm of (21b) as follows:
@@
 Now we simplify (20b) as follows:
-$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{j}(p(j|Y_{k})W(j)\right)x_{k}$
-$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k+1})-\sum_{j}(p(j|Y_{k})W(j)\right)x_{k}$ (22a)
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}(p(s^{\prime}|Y_{k})W(s^{\prime})\right)x_{k}$
+$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}(p(s^{\prime}|Y_{k})W(s^{\prime})\right)x_{k}$ (23a)
 $\;\;\;+\sum_{k=n_{0}+1}^{m}(\chi(m,k+1)a(k)-\chi(m,k)a(k-1))W(Y_{k})x_{k}$ (23b)
@@
 $\;\;\;+\chi(m,n_{0}+1)a(n_{0})W(Y_{n_{0}})x_{n_{0}}-\chi(m,m+1)a(m)W(Y_{m+1})x_{m}.$ (23d)
-Similar to the sequence $\widetilde{U}_{k+1}$, for (22a), define $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{j}p(j|Y_{k})W(j)$ for $k\geq n_{0}$ and 00 otherwise.
+Similar to the sequence $\widetilde{U}_{k+1}$, for (23a), define $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})W(s^{\prime})$ for $k\geq n_{0}$ and 0 otherwise.
 Note that $\widetilde{W}_{k+1}x_{k}$ is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
@@
 $\leq a(k)\left(K_{1}+K_{2}\|x_{k-1}\|\right)$
-$\leq a(n_{0})\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right)$
-Before simplifying (22c), we first need to repeat an important simplification from our main proof. Note that for any $0<k\leq m$,
+$\leq a(n_{0})\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right),$
+for appropriate $K_{1}$ and $K_{2}$. Before simplifying (23c), we first need to repeat an important simplification from our main proof. Note that for any $0<k\leq m$,
 $\chi(m,k)+\chi(m,k+1)a(k)=\chi(m,k+1),$
@@
 This implies that
-[…]
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\ \leq\ 1.$
 We can finally bound the norm of (23c):
@@
 Combining the bounds above gives us
-$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\widetilde{U}_{k+1}+\widetilde{W}_{k+1}x_{k}\right)+\mu_{m}(n_{0}),$
+$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right)$
+$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\widetilde{U}_{k+1}+\widetilde{W}_{k+1}x_{k}\right)+\mu_{m}(n_{0}),$
 where
-$\mu_{m}(n_{0})\leq 4a(n_{0})U_{max}+a(n_{0})W_{max}\left(K_{1}+(4+K_{2})\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right).$
+$\|\mu_{m}(n_{0})\|\leq 4a(n_{0})U_{max}+a(n_{0})W_{max}\left(K_{1}+(4+K_{2})\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right).$
 Define constants $c_{1}\coloneqq W_{max}(4+K_{2})$ and $c_{2}\coloneqq 4U_{max}+K_{1}W_{max}+c_{1}\|x^{*}\|$. This completes the proof for Lemma 2. ∎
@@
 Proof.
-We first note that, for $n_{0}\leq k\leq m$, $\|\bar{x}_{k,m}\|\leq\|\bar{x}_{k,m}-\bar{z}_{k,m}\|+\|\bar{z}_{k,m}\|\leq\bar{x}^{\prime}_{m}+\|\bar{z}_{k,m}\|$. By definition,
-$\bar{x}^{\prime}_{m}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
+We first note that, for $n_{0}\leq k\leq m$, $\|\bar{x}_{k,m}\|\leq\|\bar{x}_{k,m}-\bar{z}_{k,m}\|+\|\bar{z}_{k,m}\|\leq\bar{x}^{\prime}_{m,m}+\|\bar{z}_{k,m}\|$.
+The following follow from the definition of $\mathbb{B}_{m}$.
+If $\xi\in\mathbb{B}_{m}$, $\bar{x}^{\prime}_{m,m}(\xi)=0$ and if $\xi\notin\mathbb{B}_{m}$, $\bar{x}^{\prime}_{m,m}(\xi)=x_{m}^{\prime}(\xi)\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$.
+Hence
+$\bar{x}^{\prime}_{m,m}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
 Using (5), we have $\|\bar{z}_{k,m}\|\leq\epsilon+\|x^{*}\|$. Under the condition that $\epsilon\leq 1$ and $\delta\leq 1$, we have
-$\|\bar{x}_{k,m}\|\leq 1+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}.$
-Let $v^{(\ell)}$ denote the $\ell$th component of a vector $v$. Then
+$\|\bar{x}_{k,m}\|\leq 1+\|x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1})+1}{1-\alpha-a(n_{0})c_{1}}.$
+[…]
 $\Gamma_{m}\coloneqq$ $\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)\right\|$
@@
 Recall that $d$ here is the dimension of our iterates $\{x_{n}\}$. We apply Theorem 2 from Appendix A componentwise. For this, first note that
-$\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)^{(\ell)}\leq c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right),$
+$\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)^{(\ell)}\leq c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1})+1}{1-\alpha-a(n_{0})c_{1}}\right),$
 where $c_{3}=\max\{M_{max}+2W_{max},2U_{max}\}$. In the theorem statement, let
-$C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right),\zeta_{k,m}=\chi(m,k+1)a(k),\varepsilon=1,\gamma_{1}=1.$
+$C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1})+1}{1-\alpha-a(n_{0})c_{1}}\right),\zeta_{k,m}=\chi(m,k+1)a(k),\varepsilon=1,\gamma_{1}=1.$
 Next, we choose suitable $\gamma_{2}$ and $\omega(m)$ such that $\max_{n_{0}\leq k\leq m}\zeta_{k,m}\leq\gamma_{2}\omega(m)$.
-For this, we use our assumption that $\frac{d_{1}}{n}\leq a(n)\leq d_{3}\left(\frac{1}{n}\right)^{d_{2}},\forall\ n\geq n_{0}$, to obtain:
-$\chi(m,k+1)=\prod_{i=k+1}^{m}(1-a(i))\leq\exp\left(-\sum_{i=k+1}^{m}a(i)\right)\leq\exp\left(-\sum_{i=k+1}^{m}\frac{d_{1}}{i}\right)$
-[…]
-$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\left(\frac{k+1}{k}\right)^{d_{1}}$
-$\implies\max_{n_{0}\leq k\leq m}a(k)\chi(m,k+1)\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{k+1}{m}\right)^{d_{1}}\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{2k}{m}\right)^{d_{1}}.$
+For this, we use our assumption that $\frac{d_{1}}{n+1}\leq a(n)\leq d_{3}\left(\frac{1}{n+1}\right)^{d_{2}},\forall\ n\geq n_{0}$, to obtain:
+$\chi(m,k+1)=\prod_{i=k+1}^{m}(1-a(i))\leq\exp\left(-\sum_{i=k+1}^{m}a(i)\right)\leq\exp\left(-\sum_{i=k+1}^{m}\frac{d_{1}}{i+1}\right)$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq\exp\left(-\int_{k+1}^{m+1}\frac{d_{1}}{y+1}dy\right)\leq\exp\left(d_{1}(\log(k+2)-\log(m+2))\right)$
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\left(\frac{k+2}{m+2}\right)^{d_{1}}$
+$\implies\max_{n_{0}\leq k\leq m}a(k)\chi(m,k+1)\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{k+2}{m+2}\right)^{d_{1}}\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{2k}{m+2}\right)^{d_{1}}.$
 From the last inequality, $\gamma_{2}=d_{3}2^{d_{1}}$ and $\omega(m)=\beta_{n_{0}}(m)$ satisfy the required conditions.
@@
 $P(\Gamma_{m}\geq\delta)\leq 2de^{-D\delta/\beta_{n_{0}}(m)}.$
-The factor $d$ comes from the application of union bound to bound the maximum over all dimensions.
-Under the assumption that $\delta\leq 1$, we have that $e^{-D\delta^{2}/\beta_{n_{0}}(m)}\geq e^{-D\delta/\beta_{n_{0}}(m)}$ and hence $P(\Gamma_{m}\geq\delta)\leq 2de^{D\delta^{2}/\beta_{n_{0}}(m)}$.
+The factor $d$ comes from the application of union bound to bound the maximum over all components.
+Under the assumption that $\delta\leq 1$, we have that $e^{-D\delta^{2}/\beta_{n_{0}}(m)}\geq e^{-D\delta/\beta_{n_{0}}(m)}$ and hence $P(\Gamma_{m}\geq\delta)\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)}$.
 ∎
@@
 $\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}$
-$+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right)$
+$+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right),$
 for all $m\geq n_{0}$. Now
@@
 $\leq\exp\left(-\sum_{i=n_{0}}^{m-1}(1-\alpha)a(i)\right)\leq\exp\left(-\sum_{i=n_{0}}^{m-1}\frac{(1-\alpha)d_{1}}{i+1}\right)$
-[…]
+$\leq\exp\left(-\int_{n_{0}}^{m}\frac{(1-\alpha)d_{1}}{y+1}dy\right)\leq\exp\left((1-\alpha)d_{1}(\log(k+1)-\log(m+1))\right)$
 $=\left(\frac{n_{0}+1}{m+1}\right)^{d_{1}(1-\alpha)}\leq\frac{n_{0}+1}{m+1}.$